### v2

# On Iverson’s law of similarity

Abstract
[Iverson, 2006b] proposed the law of similarity
$\xi_{s}(\lambda x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(x)$
for the sensitivity functions $\xi_{s}\,(s\in S)$. Compared to the former models, the generality of this one lies in that here $\gamma$ and $\eta$ can also depend on the variables $\lambda$ and $s$. In the literature, this model (or its special cases) is usually considered together with a given psychophysical representation (e.g. Fechnerian, subtractive, or affine). Our goal, however, is to study at first Iverson’s law of similarity on its own. We show that if certain mild assumptions are fulfilled, then $\xi$ can be written in a rather simple form containing only one-variable functions. The obtained form proves to be very useful when we assume some kind of representation.
Motivated by [Hsu and Iverson, 2016], we then study the above model assuming that the mapping $\eta$ is multiplicatively translational. First, we show how these mappings can be characterized. Later we turn to the examination of Falmagne’s power law. According to our results, the corresponding function $\xi$ then does not have a Fechnerian representation, but it does have a subtractive representation. We close the paper with the study of the shift invariance property.
1 Introduction
[Iverson, 2006b] proposed the similarity model
$\xi_{s}(\lambda x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(x)\,,$ (1)
which we call Iverson’s law of similarity. As described shortly, this model may be applied to a number of phenomena in psychophysics and experimental psychology. The meaning of the notation in the model depends on the particular experimental context, examples of which follow. The purpose of the present work is to further the study of Iverson’s law of similarity on theoretical grounds. Models in psychology and other empirical sciences are often formalized by equations containing unknown functions, so-called functional equations. Due to this, many results in psychophysical theory can be obtained through applications of functional equations, see [Aczél, 1966], [Aczél, 1987], [Aczél et al., 2000], [Falmagne, 1985]. A seminal example of this is the work in [Luce and Edwards, 1958] in uncovering the possible ’sensory scales’ relating physical stimuli with the sensation strengths they elicit (see also [Iverson, 2006a]). Another illustrative example is the so-called Plateau’s midgray experiment (briefly described below) [Plateau, 1872], [Heller, 2006]. Regarding the theory of functional equations, we will primarily rely on J. Aczél’s monographs [Aczél, 1966] and [Aczél, 1987] and J.-C. Falmagne’s monograph [Falmagne, 1985]. In addition, we will use results from the articles [Aczél, 2005] and [Lundberg, 1977]. These statements (see Lemmas 1 and 2) are presented at the end of the second section.
The reader may wish to keep in mind the following experimental contexts for Iverson’s law of similarity (1) when considering the functional equation results in this paper. In one experimental context, the participant is asked to judge which of two stimuli has the greater sensory impact (i.e., is louder, is heavier, is brighter, etc.). The value $x$ is a non-negative number representing an intensity, and the value $\xi_{s}(x)$ indicates the intensity that is judged to be greater than intensity $x$ according to the measure of discriminability $s$. Depending on the experimental paradigm used, $s$ may be, for example, a probability, or a value proportional to $d^{\prime}$ in signal detection theory [Green and Swets, 1966], or some other measure of likelihood.111 In, for example, [Iverson, 2006b], $s$ has the following meaning. Let $P(x,y)$ be the probability that intensity $y$ is judged to be greater than intensity $x$. If such probabilities are assumed to follow a model such as $P(x,y)=F[u(y)-u(x)]$, where $F$ is strictly monotonic, then fixing $P(x,y)=\pi$, we can write that model as $F^{-1}(\pi)=u(y)-u(x)$. But note that with $\pi$ fixed, we have that $y$ in this equation depends on both $x$ and $\pi$, so we write $y=\tilde{\xi_{\pi}}(x)$, that is, $\tilde{\xi_{\pi}}(x)$ is the intensity judged greater than $x$ with probability $\pi$. It is convenient to define $F^{-1}(\pi)=s$ and also $\tilde{\xi_{\pi}}(x)=\xi_{s}(x)$. The value $\lambda$ is a non-negative real number, and the subscript $\eta(\lambda,s)$, like $s$, measures the discriminability between stimuli; the function $\eta$ may depend on $s$ or $\lambda$ (or both). The multiplier $\gamma(\lambda,s)$ similarly may depend on one or both of $s$ and $\lambda$. In this experimental context, Iverson’s law of similarity generalizes ‘Weber’s law’, for which the equation
$\xi_{s}(\lambda x)=\lambda\,\xi_{s}(x)$
holds ([Luce and Galanter, 1963]), and it also generalizes the ‘Falmagne’s power law’, for which
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{s}(x)$
holds. (We use the label ‘Falmagne’s power law’ following [Iverson, 2006b].) Falmagne’s power law gives the so-called ‘near-miss to Weber’s law’ ([McGill and Goldberg, 1968]) when the exponent $\phi(s)$ is close to but different from $1$ for some or all values of $s$ (see also [Doble et al., 2006, Augustin, 2008]). The near-miss to Weber’s law model has been applied to data from a number of experimental situations, including pure-tone intensity discrimination [Jesteadt et al., 1977], [Osman et al., 1980], [Florentine, 1986], [Florentine et al., 1987], [Doble et al., 2006], and many others), visual area discrimination ([Augustin and Roscher, 2008]), and even sugar concentration discrimination in nectar-feeding animals ([Nachev et al., 2013]).
In another experimental context, an auditory tone is embedded in a broadband noise background. The participant’s task is to match the perceived loudness of this tone/noise pairing with the loudness of an unmasked tone (a tone presented in quiet). This is the partial masking experimental context of [Pavel and Iverson, 1981]. If we write $\xi_{s}(x)$ to represent the intensity of the unmasked tone that matches the loudness of a tone of intensity $x$ embedded in a background of intensity $s$, [Pavel and Iverson, 1981] found that the data suggest the ‘shift invariance’ relationship
$\xi_{\lambda^{\theta}s}(\lambda x)=\lambda\,\xi_{s}(x),$
where $x$ and $\lambda$ vary in respective non-negative intervals, and $\theta$ takes values in the interval ]0,1[]0,1[] 0 , 1 [. This equation is another specialization of Iverson’s law of similarity.
There are still other experimental contexts for which specializations of Iverson’s law of similarity may be applied. For example, in the midgray experiment by [Plateau, 1872] alluded to above, participants were asked to paint a gray disk midway between a given white disk and a given black disk. As described in [Falmagne, 1985] and [Heller, 2006], the data suggest the relationship $\xi_{\lambda s}(\lambda x)=\lambda\,\xi_{s}(x)$, where $\xi_{s}(x)$ is the luminance of a disk judged to be midway between a disk of luminance $x$ and another of luminance $s$, and $\lambda$ is a positive number. Further examples of the application of specializations of Iverson’s law of similarity are described in [Hsu et al., 2010] and [Hsu and Iverson, 2016].
Given the above phenomena and their associated specializations of Iverson’s law of similarity, there is the natural question of the usefulness of generalizing to equation (1); after all, if the specializations of equation (1) are appropriate models for the phenomena, why would generalizations be of interest? For the case of the near-miss to Weber’s law for pure-tone intensity discrimination, generalizing to equation (1) is of practical use because models such as
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{s}(x)$
are known to fail at low intensities ([Viemeister and Bacon, 1988], [Wakefield and Viemeister, 1990], [Schroder et al., 1994]). More general candidate models are needed, and the study of equation (1) can guide experimental work in distinguishing competing models; see [Iverson, 2006b, Page 289] in this regard. In fact, deviations from Weber’s law are so common across sensory modalities (see, for example, [Carriot et al., 2021]) that similar comments may be made for a number of other intensity discrimination contexts beside those involving auditory tones. Studying generalizations of models that have been successfully applied is a theme of the present work; for instance, in Section 3, we study a generalization of the transformation $\eta(\lambda,s)=\lambda^{-\theta}s$ appearing in the ‘shift invariance’ relationship above, namely, we study the generalization of a ‘multiplicatively translational’ $\eta$,
$\eta(\lambda\tilde{\lambda},s)=\eta(\tilde{\lambda},\eta(\lambda,s))\,.$
Previously in the study of the possible functional forms in equation (1) and its specializations, researchers assumed additional representations for the function $\xi_{s}$. This approach stems historically from the original work by [Fechner, 1948], but it also has the purpose of helping to narrow down the possible solutions to the functional equations; without further assumptions about the functions $\xi_{s}$, $\eta$, and $\gamma$, equation (1) allows for too many solutions for these functions. As examples of studies of (specializations of) equation (1) in conjunction with additional representations for $\xi_{s}$, we find [Falmagne, 1994], who studied the near-miss model $\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{s}(x)$ while also assuming a ‘subtractive representation’
$s=u(\xi_{s}(x))-w(x),$
[Iverson, 2006b], who studied equation (1) in the context of a ‘Fechnerian representation’
$s=u(\xi_{s}(x))-u(x),$
[Falmagne and Lundberg, 2000], who studied the near-miss model assuming a ‘gain-control representation’
$s=\frac{u(\xi_{s}(x))-w(x)}{\sigma(x)}$
[Iverson and Pavel, 1981], who studied the shift invariance relationship
$\xi_{\lambda^{\theta}s}(\lambda x)=\lambda\,\xi_{s}(x)$
under a similar gain-control representation; [Hsu and Iverson, 2016], who studied the model
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{\eta(\lambda,s)}(x)$
under a ‘a gain-control representation’ of the form
$s=\frac{u(\xi_{s}(x))-u(x)}{\sigma(x)},$
and [Doble and Hsu, 2020], who studied equation (1) assuming a subtractive representation. As will be seen, in the present paper we examine representations such as these (Fechnerian, subtractive, and gain-control), but we also consider equation (1) on its own, especially with assumptions about the function $\eta$.
The present work is arranged as follows. Following [Falmagne, 1985], Section 2 contains the most basic concepts such as ‘psychometric families’, ‘sensitivity functions’ and some of their properties (e.g. ‘anchored’, ‘balanced’ and ‘parallel families’). Here we also present some representations known in psychophysics and their interpretations (e.g. Fechnerian, subtractive, affine, gain-control). Our results can be found in Section 3, which we divide into two subsections. First, we study equation (1), purely from the point of view of the theory of functional equations. In Theorem 1, we show that if there is at least one $s$ for which the function $\lambda\longmapsto\eta(\lambda,s)$ is invertible, then $\xi$ can written with the aid of one-variable functions in the form $\xi_{s}(x)=\dfrac{\Phi(f(s)x)}{g(s)}$. As a supplement to this, in Theorem 2 we prove that if there exists a subset $\tilde{S}\subset S$ such that for all $s\in\tilde{S}$, the mapping $J\ni\lambda\longmapsto\eta(\lambda,s)$ is constant, then $\xi$ also has a special form, namely $\xi_{s}(x)=\kappa(s)x^{\rho(s)}.$ The form obtained for the function $\xi$ in Theorem 1 proves to be very useful when we assume some kind of representation. In line with this, we first describe those functions $\xi$ that have a certain gain-control representation (see Proposition 1), and then we examine equation (1) again in the case where $\xi$ admits a representation
$\xi_{s}(x)=u(x)+v(s)$
(see Proposition 2). This latter representation is closely related to the so-called parallel families. In Proposition 3, we provide a complete description of anchored, balanced and parallel families.
Motivated by [Hsu and Iverson, 2016], in the second part of the third section we study (1) assuming that the mapping $\eta$ is multiplicatively translational. First, we show that multiplicatively translational mappings can be written in the form
$\eta(\lambda,s)=H(\lambda\cdot H^{-1}(s))$
if certain mild conditions are met. Using this form, we examine what special form the function $\gamma$ in (1) can then take. Later we turn to the examination of Falmagne’s power law, i.e.
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{\eta(\lambda,s)}(x)$
which is an important special case of (1). According to Corollary 1,
$\xi_{s}(x)=x^{\phi(s)}F(x\cdot H^{-1}(s))$
is satisfied. After that, we show that if the function $\phi$ in the exponent is monotonic, then $\phi$ is necessarily locally constant. According to Proposition 5, the corresponding function $\xi$ then does not have a Fechnerian representation, but it does have a subtractive representation (see Proposition 6). As an application of the results of the subsection, we close the paper with the study of the shift invariance
$\xi_{s}(\lambda x)=\gamma(\lambda,s)\xi_{\lambda^{\theta}s}(x).$
2 Preliminaries
Our mathematical results are applicable to each of the experimental contexts mentioned in the Introduction, but for concreteness and uniformity, we will use terminology and notation from the intensity discrimination context. (Other contexts, especially the partial masking context of [Pavel and Iverson, 1981], are referred to when especially relevant.) To this end, suppose a participant must compare a stimulus of intensity $x$ with one of intensity $y$ (both measured in ratio scale units) and judge which has the greater sensory impact (i.e., is louder, is heavier, is brighter, etc.). Let $P(x,y)$ be the probability that intensity $y$ is judged greater than intensity $x$. A simple model for these probabilities is
$P(x,y)=F(u(y)-u(x))\,,$
in which $u$ and $F$ are continuous and strictly increasing functions. This model expresses that the stimulus intensities $x$ and $y$ are scaled by the sensory mechanism to the values $u(x)$ and $u(y)$, resp., and the probabilities $P(x,y)$ are determined by the differences $u(y)-u(x)$. Equation $(\mathscr{F})$ is called a Fechnerian representation, and $u$ is termed to be a scale. The model is closely related to the problem of [Fechner, 1948]; for more details, see especially [Dzhafarov and Colonius, 1999] and [Falmagne, 1985].
This representation has been extended and generalized in several ways. There may be asymmetries between the stimuli, e.g. biases based on the order or position of stimulus presentations can occur. In this case, it is more appropriate to consider the following model
$P(x,y)=F(u(y)-w(x)),$
that is called a subtractive representation.
A still more general representation is
$P(x,y)=F(u(y)+h(y)g(x))\;\;\;\text{or}\;\;\;P(x,y)=F(u(y)h(x)+g(x)),$
which is called in the literature an affine representation; see [Hsu et al., 2010]. Affine representations include so-called gain-control representations, one of which we examine in this paper, namely, the representation
$P(x,y)=F\left(\frac{u(y)-u(x)}{\sigma(x)}\right)\,.$
This representation allows for the fact that the sensory mechanism may adjust its “gain” via the normalizing factor $\sigma(x)$. See [Hsu et al., 2010] for further description.
For all fixed $x$, introducing the function $p_{x}$ as
$p_{x}(y)=P(x,y),$
a family of functions =$\mathcal{F}=\left\{p_{x}\,|x\in I\right\}$ arises. Due to the properties of the probability, this family has (among others) the following properties
  • for all fixed $x$, the range of the function $p_{x}$ is contained in the interval ]0,1[]0,1[] 0 , 1 [
  • for all fixed $x$, the function $p_{x}$ is strictly increasing and continuous.
In what follows, our goal is to determine such families of functions under certain conditions. Therefore, following [Falmagne, 1985], we introduce the following notion.
A family of functions =$\mathcal{F}=\left\{p_{x}\,|x\in I\right\}$ is called a psychometric family, if
  1. (i)
    for all fixed $x$, the domain of the function $p_{x}$ is an open interval [^2] In this work every interval has positive length.
  2. (ii)
    for all fixed $x$, the range of the function $p_{x}$ is contained in the interval ]0,1[]0,1[] 0 , 1 [
  3. (iii)
    for all fixed $x$, the function $p_{x}$ is strictly increasing and continuous.
As we will see, instead of psychometric families, it will be much more convenient to work with sensitivity functions.
Let =$\mathcal{F}=\left\{p_{a}\,|\,a\in I\right\}$ be a psychometric family. The sensitivity function of $\mathcal{F}$ is a function $\xi$ defined for all backgrounds (or standards) $a$, and all probabilities $\pi$ in the range of a psychometric function $p_{a}$, by the equation
$\xi_{\pi}(a)=p_{a}^{-1}(\pi),$ (1)
in other words,
$p_{a}\left(\xi_{\pi}(a)\right)=\pi.$
In words: $\xi_{\pi}(a)$ is the intensity of the stimulus yielding a response probability $\pi$, for the background $a$. As J. C. Falmagne in [Falmagne, 1985] writes: ,,The change of notation, from $p_{a}^{-1}(\pi)$ to $\xi_{\pi}(a)$, symbolizes an important shift of focus in our analysis. The quantity $\pi$, the probability of the response, ceases to be the variable of interest and becomes the parameter. Typically, at most a couple of values of $\pi$ are considered in experimental plots of Weber functions or sensitivity functions. By contrast, the effect on $\xi_{\pi}(a)$ of the variable $a$ is investigated in minute detail. This is in line with a long tradition in psychophysical research in which the sensory scales uncovered by the analysis of the data are deemed of central importance.”
In terms of the sensitivity functions, $(\mathscr{F})$, $(\mathscr{S})$, and $(\mathscr{G})$ resp. can be re-formulated as
$s=u(\xi_{s}(x))-u(x),$
$s=u(\xi_{s}(x))-w(x)$
and
$s\sigma(x)=u(\xi_{s}(x))-u(x)$
We note that in each of these re-formulations, the switch from $\pi$ to $s$ occurs through the process described in footnote 1. We will use the notation with $s$ in the sequel.
If we do not assume anything about the sensitivity functions, then an infinite number of scale functions can be considered. In this paper we consider in addition a similarity law initiated by G. J. Iverson in [Iverson, 2006b]. In this case, we assume that the sensitivities fulfill identity
$\xi_{s}(\lambda x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(x)$
for all possible values of the variables $\lambda,x$ and $s$ and with appropriate two-variable functions $\gamma$ and $\eta$.
In accordance with [Falmagne, 1985], a psychometric family =$\mathcal{F}=\left\{p_{a}\,|\,a\in I\right\}$ is called anchored if there exists a number $\alpha\in]0,1[$ such that
  1. (i)
    for all $a\in I$, there exists an $x$ satisfying $p_{a}(x)=\alpha$;
  2. (ii)
    for all $x$, there exists an $a\in I$ such that $p_{a}(x)=\alpha$.
The above conditions mean in words that for every background $a$ there is a stimulus $x$, and for every stimulus $x$ there is a background $a$, such that $p_{a}(x)=\alpha$. A number $\alpha\in]0,1[$ satisfying these conditions will be called an anchor of $\mathcal{F}$.
A psychometric family $\mathcal{F}$ is called parallel if for all $p_{a},p_{b}\in\mathcal{F}$ we have
$p_{a}^{-1}(\pi)-p_{a}^{-1}(\pi^{\ast})=p_{b}^{-1}(\pi)-p_{b}^{-1}(\pi^{\ast})$
whenever all four terms are defined.
If a parallel psychometric family =$\mathcal{F}=\left\{p_{a}\,|\,a\in I\right\}$ is given, then and only then
$\xi_{\pi}(a)-\xi_{\pi^{\ast}}(a)=\xi_{\pi}(b)-\xi_{\pi^{\ast}}(b)$
is fulfilled by the sensitivity functions. In view of [Falmagne, 1985, Theorem 8.5], an anchored psychometric family =$\mathcal{F}=\left\{p_{a}\,|\,a\in I\right\}$ is parallel if and only if there exist strictly increasing and continuous functions $u$ and $v$ such that
$\xi_{s}(x)=u(x)+v(s)$
for all possible $x$ and $s$.
Intuitively, a psychometric family is parallel if any two psychometric functions can be made to coincide by a horizontal shift of one toward the other. This suggests that given one psychometric function, say $p_{a}$, any other psychometric function $p_{b}$ is completely characterized by the value of one parameter depending on $b$ that is denoted by $g(b)$, expressing the length and direction of the above-mentioned shift ($g(b)$ may be negative).
A psychometric family =$\mathcal{F}=\left\{p_{a}\,|\,a\in I\right\}$ is called balanced if
$p_{a}(b)+p_{b}(a)=1$
holds for all possible $a$ and $b$.
The assumption that the psychometric family =$\mathcal{F}=\left\{p_{a}\,|\,a\in I\right\}$ is balanced holds if and only if for the sensitivity function we have
$\xi_{1-\pi}\left(\xi_{\pi}(a)\right)=a$
for all $\pi$ and $a$.
As J. Cl. Falmagne in [Falmagne, 1985] writes, there are cases (i.e. the brightness discrimination experiment) where the balance condition naturally holds by the paradigm or of the method of data collection. At the same time, observe that the order of the stimuli in the notation $(a,b)$ suggests the fact that stimulus $a$ is presented first, followed by stimulus $b$. The balance condition states, in effect, that the order of presentation does not affect the result of a comparison. This may not be true, for example, for auditory stimuli presented successively.
These properties may seem natural due to the aforementioned, but we will see later that these notions are far too restrictive, even without Iverson’s law of similarity, see Propositions 2 and 3.
To prove Proposition 1, the lemma below will be utilized from [Aczél, 2005].
Lemma 1.
Let $D\subset\mathbb{R}^{2}$ be an open and connected set. Assume that the real-valued functions $f,g,k,h$ fulfill
$f(st)=g(s)+h(s)\cdot k(t)$
for all $(s,t)\in D$. If the function $f$ is measurable and philandering, then
  1. (A)
    either
    $\begin{array}[]{rcl}f(r)&=&\beta_{0}\ln(r)+\beta_{1}+\beta_{2}\beta_{3}\\ g(s)&=&\beta_{0}\ln(s)+\beta_{1}\\ h(s)&=&\beta_{3}\\ k(t)&=&\frac{\beta_{0}}{\beta_{3}}\ln(t)+\beta_{2}\end{array}$
  2. (B)
    or
    $\begin{array}[]{rcl}f(r)&=&\beta_{3}\beta_{4}r^{\beta_{0}}+\beta_{1}\\ g(s)&=&\beta_{2}s^{\beta_{0}}+\beta_{1}\\ h(s)&=&\beta_{3}s^{\beta_{0}}\\ k(t)&=&\beta_{4}t^{\beta_{0}}-\frac{\beta_{2}}{\beta_{3}}\end{array}$
for all $s\in\left\{s\,|\,\text{there exists $t$ such that $(s,t)\in D$}\right\}$, $t\in\left\{t\,|\,\text{there exists $s$ such that $(s,t)\in D$}\right\}$ and $r\in\left\{st\,|\,(s,t)\in D\right\}$, where $\beta_{i}$ are real constants for $i=1,2,3,4$ such that $\beta_{0},\beta_{3},\beta_{4}\neq 0$.
Further, in the next section we also use a result of [Lundberg, 1977] on the functional equation
$f(\ell(y)+g(s))=m(y)+h(y+s)$
which is the statement below.
Lemma 2.
Suppose the equation
$f(\ell(x)+g(y))=m(x)+h(x+y)$ (2)
holds for all $(x,y)$ in a body in $\mathbb{R}^{2}$, where $f$, $g$, $h$, $\ell$ and $m$ are real-valued, continuous functions, each defined on an interval (possibly different intervals for different functions). Suppose also that $h$, $\ell$ and $m$ are philandering. Then the solutions are exactly those given in Cases I-V below, where the parameters can take any real values as long as the resulting expressions are real numbers.
  1. Case I:
    $f(x)$ $=\alpha+\rho x$
    $g(x)$ $=\beta+bx$
    $h(x)$ $=-\tau+\rho bx$
    $\ell(x)$ is arbitrary
    $m(x)$ $=\rho\ell(x)-\rho bx+\alpha+\rho\beta+\tau$
  2. Case II:
    $f(x)$ $=\alpha+\rho\log(c+e^{\kappa x})$
    $g(x)$ $=\frac{1}{\kappa}\log(-\beta c+de^{\delta x})$
    $h(x)$ $=-\tau+\alpha+\rho\log(bc+de^{\delta x})$
    $\ell(x)$ $=-\frac{1}{\kappa}\log(\beta+be^{-\delta x})$
    $m(x)$ $=\tau-\rho\log(b+\beta e^{\delta x})$
  3. Case III:
    $f(x)$ $=\rho\log(\alpha-be^{\kappa x})$
    $g(x)$ $=\frac{1}{\kappa}\log(\beta-d\alpha x)$
    $h(x)$ $=-\tau+\rho\log(bd\alpha x+\alpha\epsilon-b\beta)$
    $\ell(x)$ $=-\frac{1}{\kappa}\log(\epsilon+bdx)$
    $m(x)$ $=\tau-\rho\log(\epsilon+bdx)$
  4. Case IV:
    $f(x)$ $=\alpha+\rho e^{\kappa x}$
    $g(x)$ $=\beta+\frac{1}{\kappa}\log(b+ce^{\delta x})$
    $h(x)$ $=-\tau+\alpha+\rho ce^{\delta x}$
    $\ell(x)$ $=-\beta+\frac{\delta}{\kappa}x$
    $m(x)$ $=\tau+\rho be^{\delta x}$
  5. Case V:
    $f(x)$ $=\alpha+\frac{\rho}{\delta}\log(\beta+x)$
    $g(x)$ $=-\beta-\epsilon+ce^{\delta x}$
    $h(x)$ $=-\tau+\alpha+\frac{\rho}{\delta}\log(b+ce^{\delta x})$
    $\ell(x)$ $=\epsilon+be^{-\delta x}$
    $m(x)$ $=\tau-\rho x$
We end this section with some short comments. Recall from footnote 2, that all intervals in this paper have positive length. Another is that in several results below, we assume that the value $s=0$ is attainable. Note that $s=0$ is feasible in all of the experimental contexts mentioned in the Introduction. In the case of intensity discrimination experiments, the index $s=0$ may be assumed to correspond to the probability of $1/2$ (or to the $d^{\prime}$ value of 00), and for other experiments such as partial masking, the intensity $s=0$ would also make sense for the model. The third short comment is that even though we are using terminology adopted from an intensity discrimination context, we will continue to use that terminology when referring to other experimental contexts. For example, we use the term ‘sensitivity function’ even in a partial masking context.
3 Results
3.1 Iverson’s law of similarity
According to our primary aim, now we study Iverson’s law of similarity, i.e., identity
$\xi_{s}(\lambda x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right).$ (3)
In the following statement, we show that if there exists an $s^{\ast}\in S$ such that the mapping $\lambda\longmapsto\eta(\lambda,s^{\ast})$ is invertible, then the function $\xi$ can be expressed with the aid of one-variable functions in a rather special form.
Before this, however, we would like to clarify under which conditions we will be able to prove this statement. Additionally, we would like to compare our assumptions here with the standard assumptions found in the literature. In what follows (until stated otherwise), $\xi$ will always denote a one-parameter family that is defined on the Cartesian product $I\times S$ of some real intervals $I$ and $S$. This means that we do not assume that $\xi$ is a sensitivity function, which among other things would imply that for all fixed $s$, the mapping $\xi_{s}$ is continuous and strictly increasing. Further, the domain of the two-variable function $\gamma$ and $\eta$ is supposed to be $J\times S$. In the case of the function $\gamma$, there is no assumption on the range, while in the case of the function $\eta$ we assume that it maps $J\times S$ into $S$. This latter assumption is justified by equation (3) itself. It is important to emphasize that apart from these, we do not impose any a priori regularity assumptions on the involved functions, but in several cases, we get them as a consequence. Finally, the assumption $I\cdot J\subset I$ helps equation (3) to hold for all $x\in I$, $\lambda\in J$ and $s\in S$. Thus, from one point of view, in this paper we will work under more general conditions, since in most papers the authors assume that the functions $\gamma$ and $\eta$ are continuous in each of their variables. On the other hand, the domain of $\xi$ is assumed to be $I\times S$, which, compared to other works, is a bit more special. As an example, we mention that in [Doble and Hsu, 2020], the parameters $s$ are taken from an open interval $S$, for all $s\in S$, the mapping $\xi_{s}$ is defined on an open interval $I_{s}\subset]0,+\infty[$ such that the set
$B=\left\{(x,s)\,|\,s\in S,x\in I_{s}\right\}\subset\mathbb{R}^{2}$
is an open and connected set and equation (3) is supposed to hold for all triple $(x,\lambda,s)$ such that $\lambda\in J$, $(x,s)\in B$ and $(\lambda x,s)\in B$.
Theorem 1.
Let $I,J,S\subset\mathbb{R}$ be intervals with $I\cdot J=\left\{\lambda x\,|\,x\in I,\lambda\in J\right\}\subset I$. Let further $\gamma\colon J\times S\to\mathbb{R}$ and $\eta\colon J\times S\to S$ be functions and assume that the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ fulfills
$\xi_{s}(\lambda x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right).$
If there exists an $s^{\ast}\in S$ such that the mapping $\lambda\longmapsto\eta(\lambda,s^{\ast})$ is invertible, then there exist functions $f\colon\tilde{S}\to J$, $g\colon\tilde{S}\to\mathbb{R}$ and $\Phi\colon I\to\mathbb{R}$ such that
$g(s)\cdot\xi_{s}(x)=\Phi(f(s)\cdot x)\qquad\left(x\in I,s\in\tilde{S}\right).$
Here $\tilde{S}$ denotes the range of the mapping $\lambda\longmapsto\eta(\lambda,s^{\ast})$, that is, $\tilde{S}=\eta(J,s^{\ast})$.
Proof.
Let $I,J,S\subset\mathbb{R}$ be intervals with $I\cdot J=\left\{\lambda x\,|\,x\in I,\lambda\in J\right\}\subset I$. Let further $\gamma\colon J\times S\to\mathbb{R}$ and $\eta\colon J\times S\to S$ be functions and assume that the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ fulfills
$\xi_{s}(\lambda x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right).$
By our assumptions, there exists an $s^{\ast}\in J$ such that the mapping $\lambda\longmapsto\eta(\lambda,s^{\ast})$ is invertible. Then the inverse of this mapping is defined on the range of $\lambda\longmapsto\eta(\lambda,s^{\ast})$, that is, on $\tilde{S}=\eta(J,s^{\ast})$. Therefore, the mapping
$J\ni\lambda\longmapsto\eta(\lambda,s^{\ast})\in\tilde{S}$
is a bijection between the sets $J$ and $\tilde{S}$. In other words, for all $s\in\tilde{S}$, there exists a uniquely determined $\lambda\in J$ such that $\eta(\lambda,s^{\ast})=s$, or equivalently $\lambda=\eta^{-1}(s,s^{\ast})$.
The above identity with $s^{\ast}$ instead of $s$ is just
$\xi_{s^{\ast}}(\lambda x)=\gamma(\lambda,s^{\ast})\xi_{\eta(\lambda,s^{\ast})}(x)\qquad\left(x\in I,\lambda\in J\right).$
Let now $s\in\tilde{S}$ be arbitrary. From this, with the substitution $s=\eta^{-1}(\lambda,s^{\ast})$ we get that
$\xi_{s^{\ast}}(\eta^{-1}(s,s^{\ast})x)=\gamma(\eta^{-1}(s,s^{\ast}),s^{\ast})\xi_{s}(x)\qquad\left(x\in I,s\in\tilde{S}\right).$
Let now
$f(s)=\eta^{-1}(s,s^{\ast})\qquad g(s)=\gamma(\eta^{-1}(s,s^{\ast}),s^{\ast})\qquad\text{and}\qquad\Phi(x)=\xi_{s^{*}}(x)\qquad\left(x\in I,s\in\tilde{S}\right)$
to deduce the statement. ∎
Remark 1.
Using the notations and the assumptions of Theorem 1, if the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in J)$ fulfills
$\xi_{s}(\lambda x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,\lambda\in J\setminus\left\{0\right\},s\in\tilde{S}\right),$
then $\gamma(\lambda^{\ast},s)=0$ implies that $\xi_{s}(x)=0$ on some subinterval of $I$. Indeed, let $s\in\tilde{S}$ be fixed and assume that there exists $\lambda^{\ast}\in J\setminus\left\{0\right\}$ such that $\gamma(\lambda^{\ast},s)=0$, then
$\xi_{s}(\lambda^{\ast}x)=\gamma(\lambda^{\ast},s)\xi_{\eta(\lambda^{\ast},s)}\left(x\right)=0$
for all $x\in I$, yielding that $\xi_{s}$ is identically zero on the interval $\lambda^{\ast}I=\left\{\lambda^{\ast}x\,|\,x\in I\right\}$.
In some contexts, it is assumed that $\xi_{s}(x)>0$ for all possible $s$ and all possible $x$. Note that in such a case it follows from Theorem 1, that we have
$\xi_{s}(x)=\frac{\Phi(f(s)x)}{g(s)}$
for all $x\in I$ and $s\in\tilde{S}$.
Remark 2.
The assumption that there exists $s^{\ast}\in S$ such that the mapping
$J\ni\lambda\longmapsto\eta(\lambda,s^{\ast})$
is invertible, is essential while proving Theorem 1. However, from the point of view of applications, this condition may seem artificial. Instead of invertibility, for example, monotonicity is a much more reasonable assumption. Note, however, that then the following alternative is fulfilled. Accordingly, suppose that there exists an $s^{\ast}\in S$ such that the mapping $\lambda\longmapsto\eta(\lambda,s^{\ast})$ is monotonic on the interval $J$. Then the points of discontinuity of this function form a countable set. These points of discontinuity induce a (countable) disjoint partition $(J_{\alpha})_{\alpha\in A}$ of the interval $J$. Then for all $\alpha\in A$ the mapping
$J^{\circ}_{\alpha}\ni\lambda\longmapsto\eta(\lambda,s^{\ast})$
is continuous (and monotonic), where $J^{\circ}_{\alpha}$ denotes the interior of the interval $J_{\alpha}$. Then there are subintervals $J_{\alpha,c},J_{\alpha,s}\subset J^{\circ}_{\alpha}$ such that the above mapping is constant on $J_{\alpha,c}$ and strictly monotonic on $J_{\alpha,s}$. The mapping
$\lambda\longmapsto\eta(\lambda,s^{\ast})$
while restricted to $J_{\alpha,s}$ is invertible. Thus the assumptions of Theorem 1 are fulfilled on $J_{\alpha,s}$. Further the same mapping, while restricted to $J_{\alpha,c}$, is a constant mapping. This motivates the following theorem, in which Iverson’s law of similarity is examined in the case where the mapping $\eta$ is constant for certain values $s$. Notice that in this statement, in contrast to Theorem 1, we only need to require the minimum about the domains of the involved functions.
Theorem 2.
Let $S$ and $J$ be intervals and suppose that for all $s\in S$, there exists an interval $I_{s}\subset]0,+\infty[$ such that the set
$B=\left\{(x,s)\,|\,s\in S,x\in I_{s}\right\}\subset\mathbb{R}^{2}$
is open and connected. Suppose further that
  1. (i)
    we have
    $\xi_{s}(x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(x)$
    for all $\lambda\in J$ and $(x,s)\in B$ for which also $(\lambda x,s)\in B$,
  2. (ii)
    there exists a subset $\tilde{S}\subset S$ such that for all $s\in\tilde{S}$, the mapping
    $J\ni\lambda\longmapsto\eta(\lambda,s)$
    is constant,
  3. (iii)
    for all $s\in\tilde{S}$, the mapping
    $I_{s}\ni x\longmapsto\xi_{s}(x)$
    is measurable.
Then there exists functions $\kappa,\rho\colon\tilde{S}\to\mathbb{R}$ such that
$\xi_{s}(x)=\kappa(s)x^{\rho(s)}$
holds for all $s\in\tilde{S}$ and $x\in I_{s}$.
Proof.
Under the assumptions of the theorem, let $s^{\ast}\in\tilde{S}$ be arbitrary. Then, the mapping
$J\ni\lambda\longmapsto\eta(\lambda,s^{\ast})$
is constant. Thus Iverson’s law of similarity with $s=s^{\ast}$, i.e.
$\xi_{s^{\ast}}(x)=\gamma(\lambda,s^{\ast})\xi_{\eta(\lambda,s^{\ast})}(x)\qquad\left(x\in I_{s},\lambda\in J\right),$
becomes a multiplicative Pexider equation. Further, the function $\xi_{s^{\ast}}$ is measurable on the interval $I_{s^{\ast}}$. Thus, there exist real constants (depending on $s^{\ast}$), say $\kappa(s^{\ast})$ and $\rho(s^{\ast})$ such that
$\xi_{s^{\ast}}(x)=\kappa(s^{\ast})x^{\rho(s^{\ast})}\qquad\left(x\in I_{s^{\ast}}\right).$
Since $s^{\ast}\in\tilde{S}$ was an arbitrary element of the set $\tilde{S}$, this procedure can be repeated for any element of the set $\tilde{S}$. So, finally, we obtain that there exist functions $\kappa,\rho\colon\tilde{S}\to\mathbb{R}$ such that
$\xi_{s}(x)=\kappa(s)x^{\rho(s)}$
for all $s\in\tilde{S}$ and $x\in I_{s}$. ∎
Remark 3.
It can be seen from the steps of the above proof that we do not necessarily have to assume that the function $\xi_{s}$ is measurable for all $s\in\tilde{S}$. However, then we only get a ‘local’ version of the above statement. More precisely, if we assume that there is an $s^{\ast}$ for which $\xi_{s^{\ast}}$ is measurable, then we get that there exist real constants $\kappa(s^{\ast})$ and $\rho(s^{\ast})$ such that
$\xi_{s^{\ast}}(x)=\kappa(s^{\ast})x^{\rho(s^{\ast})}\qquad\left(x\in I_{s^{\ast}}\right).$
Remark 4.
Notice that the representation obtained in the above theorem satisfies an identity much more specific than Iverson’s law of similarity. Indeed, if
$\xi_{s}(x)=\kappa(s)x^{\rho(s)}$
is fulfilled for all $s\in\tilde{S}$ and $x\in I_{s}$, then
$\xi_{s}(\lambda x)$ $=\kappa(s)(\lambda x)^{\rho(s)}$
$=\kappa(s)\lambda^{\rho(s)}x^{\rho(s)}$
$=\lambda^{\rho(s)}\cdot\kappa(s)x^{\rho(s)}$
$=\lambda^{\rho(s)}\xi_{s}(x)$
for all $\lambda\in J$ and $(x,s)\in B$ for which also $(\lambda x,s)\in B$, that is, Falmagne’s power law model holds. As mentioned, in the case of intensity discrimination, this model is called the near-miss to Weber’s law when the exponent $\rho(s)$ is close to but different from $1$ for some or all $s$, and it gives Weber’s law when $\rho(s)=1$ for all $s$.
So we see from Theorems 1 and 2 that assumptions about $\eta$ give helpful information about Iverson’s law of similarity without assuming a further representation for $\xi$. Theorem 2 says that for $s$ for which the mapping $\lambda\longmapsto\eta(\lambda,s)$ is constant, Iverson’s law of similarity simplifies to the Falmagne’s power law model. In the extreme case that this mapping is constant for all $s$ (that is, $\eta(\lambda,s)=\nu(s)$ for some function $\nu$, for all $(\lambda,s)$), Falmagne’s power law model holds for all $s$. Theorem 1 looks at a case for which there is some $s^{\ast}$ for which this does not hold, and more specifically, that there is some $s^{\ast}$ such that $\lambda\longmapsto\eta(\lambda,s^{\ast})$ is invertible. In such a case, Theorem 1 shows that the function $\xi$ can be expressed with the aid of one-variable functions.
In the following, we determine those mappings which have the form that appears in Theorem 1 and that also have a gain-control type affine representation of the form
$\xi_{s}(x)=u^{-1}\left(s\sigma(x)+u(x)\right)$
for all possible $x$ and $s$, and with appropriate functions $u$ and $\sigma$.
Proposition 1.
Let $I,J,S\subset\mathbb{R}$ be intervals and suppose that the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ admits a representation
$\xi_{s}(x)=\frac{\Phi(f(s)x)}{g(s)}\qquad\left(x\in I,s\in S\right)$
with some functions $\Phi,f$ and $g$. Assume further, that the mapping $\xi$ fulfills the following affine representation
$\xi_{s}(x)=u^{-1}\left(s\sigma(x)+u(x)\right)$
for all possible $x\in I$ and $s\in S$ with an appropriate strictly monotonic function $u$ and with an appropriate function $\sigma$. If $0\in S$ and $g(0)\neq 0$, then
  1. (A)
    either the functions $u$ and $\sigma$ are affine transformations of the same multiplicative function, more precisely,
[…]
    and we also have
    $\xi_{s}(x)=\sqrt[\mu]{\frac{\alpha_{2}}{\alpha_{1}}s+1}\cdot x\qquad\left(x\in I,s\in S\right)$
  2. (B)
    or the function $\sigma$ is constant and $u$ is an affine transform of the logarithm function, more precisely,
[…]
    and therefore
    $\xi_{s}(x)=\exp\left(\frac{s\sigma}{\beta_{0}}\right)\cdot x\qquad\left(x\in I,s\in S\right)$
  3. (C)
    or the function $\sigma$ is the identically zero function, $u$ is an arbitrary strictly monotonic function and we have
    $\xi_{s}(x)=x\qquad\left(x\in I,s\in S\right).$
Proof.
Under the assumptions of the proposition, we have
$\frac{\Phi(f(s)x)}{g(s)}=u^{-1}\left(s\sigma(x)+u(x)\right)\qquad\left(x\in I,s\in S\right).$
Let $s=0$ in this identity to deduce that
$\frac{\Phi(f(0)x)}{g(0)}=u^{-1}(u(x))=x\qquad\left(x\in I\right),$
that is,
$\Phi(x)=\alpha x\qquad\left(x\in I\right),$
where $\alpha=\dfrac{g(0)}{f(0)}$. Note that this already yields that
$\xi_{s}(x)=\alpha(s)x\qquad\left(x\in I,s\in S\right)$
with an appropriate function $\alpha\colon S\to\mathbb{R}$.
Therefore
$\alpha(s)x=u^{-1}(s\sigma(x)+u(x))\qquad\left(x\in I,s\in S\right),$
or equivalently
$u(\alpha(s)x)=s\sigma(x)+u(x)\qquad\left(x\in I,s\in S\right).$
If $\sigma$ is the identically zero function, then we obtain
$\alpha(s)x=x\qquad\left(x\in I,s\in S\right),$
from which $\alpha(s)=1$ follows for all $s\in S$. This leads to alternative (C).
If $\sigma$ is not the identically zero function, then there exists an $x^{\ast}\in I$ such that $\sigma(x^{\ast})\neq 0$. Note that if such a point $x^{\ast}$ exists, then we necessarily have $x^{\ast}\neq 0$. Indeed, if we would have $x^{\ast}=0$, then the above identity would take the following form
$u(0)=s\sigma(0)+u(0)\qquad\left(s\in S\right),$
which is impossible. Further, the above identity with $x^{\ast}$ instead of $x$ is
$\alpha(s)x^{\ast}=u^{-1}(s\sigma(x^{\ast})+u(x^{\ast}))\qquad\left(s\in S\right).$
Since the function $u$ is strictly monotonic and continuous, the right hand side, as a function of the variable $s$, i.e., the mapping
$s\longmapsto u^{-1}(s\sigma(x^{\ast})+u(x^{\ast}))$
is also strictly monotonic and continuous. From this we obtain that the function $\alpha$ is strictly monotonic and continuous. Thus the set $\alpha(J)=\left\{\alpha(s)\,|\,s\in S\right\}$ is an interval, and the function $\alpha$ is invertible. Therefore we have
$u(tx)=\alpha^{-1}(t)\sigma(x)+u(x)\qquad\left(x\in I,t\in\alpha(J)\right).$
In view of Lemma 1333To apply Lemma 1, the intervals $I$ and $S$ should be open intervals. This may not be fulfilled. At the same time, if this is the case, then we work on $I^{\circ}$, resp. on $J^{\circ}$, i.e., on the interiors of these intervals. The theorem of Aczél is then applicable to these open intervals. Note, however, that when we move to the interiors $I^{\circ}$, resp. $S^{\circ}$ of the intervals $I$ and $S$, we ”lose” at most two points, the endpoints. However, since the functions in the statement are continuous, the function values at the points in question can be obtained with taking the limits at the end of the proof., either there exist real constants $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$ and $\mu$ such that
$u(x)=\alpha_{1}x^{\mu}+\beta_{1}\qquad\text{and}\qquad\sigma(x)=\alpha_{2}x^{\mu}+\beta_{2}\qquad\left(x\in I\right)$
or there exist real constants $\sigma,\beta_{0}\neq 0$ and $\beta_{1}$ such that
$u(x)=\beta_{0}\ln(x)+\beta_{1}\qquad\text{and}\qquad\sigma(x)=\sigma\qquad\left(x\in I\right).$
In the first case, we find that necessarily $\beta_{2}=0$ so
$\xi_{s}(x)=\sqrt[\mu]{\frac{\alpha_{2}}{\alpha_{1}}s+1}\cdot x\qquad\left(x\in I,s\in S\right),$
while in the second case, we have
$\xi_{s}(x)=\exp\left(\frac{s\sigma}{\beta_{0}}\right)\cdot x\qquad\left(x\in I,s\in S\right).$
Remark 5.
If the conditions of the above proposition are fulfilled, then we deduce that in all of alternatives (A)–(C), the mapping $\xi_{s}$ must satisfy Weber’s law, i.e. $\xi_{s}(\lambda x)=\lambda\xi_{s}(x)$ must hold for all possible values of $x$ and $s$. As Weber’s law is too restrictive for many experimental situations, we conclude that the assumptions of the proposition (the forms $\xi_{s}(x)=\frac{\Phi(f(s)x)}{g(s)}$ and $\xi_{s}(x)=u^{-1}\left(s\sigma(x)+u(x)\right)$, and the fact that $0\in S$) are quite limiting. (Especially, we see that Alternative (C), for which $\xi_{s}$ does not depend on $s$, would not occur in any applications under consideration.) However, see Remark 6 and the representation considered in Proposition 2 below.
Remark 6.
Let $I,S=]0,+\infty[$ and consider the sensitivity functions $\xi_{s}$ defined on $I$ for all $s\in S$ by
$\xi_{s}(x)=x+s\qquad\left(x\in I,s\in S\right).$
These functions fulfill Iverson’s law of similarity since we have
$\xi_{s}(\lambda x)=s+\lambda x=\lambda\left(\frac{s}{\lambda}+x\right)=\lambda\xi_{\frac{s}{\lambda}}(x)$
for all $\lambda,x\in I$ and $s\in S$, that is, in this case
$\gamma(\lambda,s)=\lambda\qquad\text{and}\qquad\eta(\lambda,s)=\frac{s}{\lambda}\qquad\left(\lambda\in I,s\in S\right).$
Moreover, in accordance with Theorem 1 we have
$\xi_{s}=\frac{\Phi(f(s)x)}{g(s)}\qquad\left(x\in I,s\in S\right),$
since
$\xi_{s}(x)=s+x=s\left(1+\frac{x}{s}\right)=\dfrac{\left(1+\frac{x}{s}\right)}{\frac{1}{s}}=\dfrac{\left(1+\frac{1}{s}x\right)}{\frac{1}{s}}\qquad\left(x\in I,s\in S\right),$
i.e.
$f(s)=\frac{1}{s}\qquad g(s)=\frac{1}{s}\qquad\text{and}\qquad\Phi(x)=x+1\qquad\left(x\in I,s\in S\right).$
Finally, we have
$\xi_{s}(x)=u^{-1}(s+u(x))\qquad\left(x\in I,s\in S\right),$
i.e. a gain-control type affine representation with
$u(x)=x\qquad\text{and}\qquad\sigma(x)=1\qquad\left(x\in I\right).$
Nevertheless, the sensitivity functions $\xi_{s}$ cannot be written in any of the forms (A)–(C) in Proposition 1. This is caused by the fact that in this example the assumption $0\in S$ does not hold.
We now examine the form $\xi_{s}(x)=\frac{\Phi(f(s)x)}{g(s)}$ along with a different representation than in Proposition 1, namely, we examine the representation $\xi_{s}(x)=u(x)+v(s)$. Recall that this is the representation associated with an anchored and parallel psychometric family, although here we do not assume that $u$ and $v$ are continuous.
Proposition 2.
Let $I,S\subset\mathbb{R}$ be intervals and suppose that the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in J)$ admits a representation
$\xi_{s}(x)=\frac{\Phi(f(s)x)}{g(s)}\qquad\left(x\in I,s\in S\right)$
with some functions $\Phi,f$ and $g$. Suppose further that the mapping $\xi$ admits a representation of the form
$\xi_{s}(x)=u(x)+v(s)\qquad\left(x\in I,s\in S\right).$ (3)
If the function $\Phi$ is measurable and philandering and $f$ is strictly monotonic and continuous, then
  1. (a)
    either there exist real constants $\alpha,\beta,\gamma$ with $\alpha\neq 0$ and a function $f\colon S\to\mathbb{R}$ such that
    $u(x)=\alpha\ln(x)+\beta\qquad v(s)=\alpha\ln(s)+\gamma\qquad\left(x\in I,s\in S\right)$
    and
    $\xi_{s}(x)=\alpha\ln(f(s)x)+\beta+\gamma\qquad\left(x\in I,s\in S\right)$
  2. (b)
    or there exist real constants $\alpha,\beta,\gamma,\rho$ with $\alpha,\beta,\rho\neq 0$ and a non-zero function $f\colon S\to\mathbb{R}$ such that
    $u(x)=\alpha x^{\rho}-\beta\qquad v(s)=\frac{\gamma}{f(s)^{\rho}}+\beta\qquad\left(x\in I,s\in S\right)$
    and
    $\xi_{s}(x)=\alpha x^{\rho}+\frac{\gamma}{f(s)^{\rho}}\qquad\qquad\left(x\in I,s\in S\right).$
Proof.
Under the assumptions of the proposition, there exist functions $f,g$ defined on $S$ and a function $\Phi$ defined on $I$ such that
$g(s)\cdot\xi_{s}(x)=\Phi(f(s)\cdot x)\qquad\left(x\in I,s\in S\right).$
On the other hand, due to the other representation, we have
$g(s)\xi_{s}(x)=g(s)u(x)+g(s)v(s)$
for all $x\in I$ and $s\in S$. Thus
$\Phi(f(s)\cdot x)=g(s)u(x)+g(s)v(s)\qquad\left(x\in I,s\in S\right).$
Since the function $f$ is strictly monotonic and continuous, the set $f(S)$ is an interval and $f$ is invertible. Thus we get that
$\Phi(sx)=g(f^{-1}(s))u(x)+g(f^{-1}(s))v(f^{-1}(s))$
for all $s\in f(S)$ and $x\in I$. Further, the function $\Phi$ as assumed to be measurable and philandering, therefore Lemma 1 applies. This means that
  1. (a)
    either
    $\begin{array}[]{rcl}\Phi(x)&=&\beta_{0}\ln(x)+\beta_{1}+\beta_{2}\beta_{3}\\ g(f^{-1}(s))v(f^{-1}(s))&=&\beta_{1}\ln(s)+\beta_{1}\\ g(f^{-1}(s))&=&\beta_{3}\\ u(x)&=&\frac{\beta_{0}}{\beta_{3}}\ln(x)+\beta_{2}\end{array}$
  2. (b)
    or
    $\begin{array}[]{rcl}\Phi(x)&=&\beta_{3}\beta_{4}x^{\beta_{0}}+\beta_{1}\\ g(f^{-1}(s))v(f^{-1}(s))&=&\beta_{2}s^{\beta_{0}}+\beta_{1}\\ g(f^{-1}(s))&=&\beta_{3}s^{\beta_{0}}\\ u(x)&=&\beta_{4}x^{\beta_{0}}-\frac{\beta_{2}}{\beta_{3}},\end{array}$
    $\beta_{i}$ are real constants for $i=1,2,3,4$ such that $\beta_{0},\beta_{3},\beta_{4}\neq 0$.
Let us consider alternative (a) first. The third identity shows that the function $g$ is constant. Using this in the second identity,
$\beta_{3}v(f^{-1}(s))=\beta_{1}\ln(s)+\beta_{1},$
i.e.,
$v(s)=\frac{\beta_{1}}{\beta_{3}}\ln(f(s))+\frac{\beta_{1}}{\beta_{3}}$
follows for all $s\in S$. So
$\xi_{s}(x)=u(x)+v(s)=\frac{\beta_{0}}{\beta_{3}}\ln(x)+\beta_{2}+\frac{\beta_{1}}{\beta_{3}}\ln(f(s))+\frac{\beta_{1}}{\beta_{3}}.$
At the same time, we also have to have
$g(s)\xi_{s}(x)=\Phi(f(s)x).$
Comparing the coefficients, this is possible only if $\beta_{0}=\beta_{1}$. Therefore
$\begin{array}[]{rcl}u(x)&=&{\dfrac{{\beta}_{1}\,\ln(x)}{{\beta}_{3}}}+{\beta}_{2}\\[8.53581pt] v(s)&=&{\dfrac{{\beta}_{1}\,\ln(f\left(s\right))}{{\beta}_{3}}}+{\dfrac{{\beta}_{1}}{{\beta}_{3}}}\end{array}\qquad\left(x\in I,s\in S\right)$
and
$\xi_{s}(x)={\frac{{\beta}_{1}\,\ln(x)}{{\beta}_{3}}}+{\frac{{\beta}_{1}\,\ln(f\left(s\right))}{{\beta}_{3}}}+{\frac{{\beta}_{1}}{{\beta}_{3}}}+{\beta}_{2}$
for all $x\in I$ and $s\in S$.
Finally, consider alternative (b). The third identity yields that
$g(s)=\beta_{3}f(s)^{\beta_{0}}\qquad\left(s\in S\right)$
which together with the second identity implies that
$v(x)={\frac{{\beta}_{2}\,f\left(s\right)^{{\beta}_{0}}+{\beta}_{1}}{{\beta}_{3}\,f\left(s\right)^{{\beta}_{0}}}}=\frac{\beta_{2}}{\beta_{3}}+\frac{\beta_{1}}{\beta_{3}}f(s)^{-\beta_{0}}.$
Since
$u(x)=\beta_{4}x^{\beta_{0}}-\frac{\beta_{2}}{\beta_{3}}\qquad\left(x\in I\right),$
we obtain that
$\xi_{s}(x)={\beta}_{4}\,x^{{\beta}_{0}}+{\frac{{\beta}_{2}\,f\left(s\right)^{{\beta}_{0}}+{\beta}_{1}}{{\beta}_{3}\,f\left(s\right)^{{\beta}_{0}}}}-{\frac{{\beta}_{2}}{{\beta}_{3}}}={\beta}_{4}\,x^{{\beta}_{0}}+\frac{\beta_{1}}{\beta_{3}}f(s)^{-\beta_{0}}\qquad\left(x\in I,s\in S\right).$
Remark 7.
Observe that in contrast to Proposition 1, in Proposition 2 there was no need to assume that $0\in S$. Thus, the sensitivity function that was considered in Remark 6, appears here as a possible solution. Indeed, let us take
$\alpha=\gamma=\rho=1\quad\text{and}\quad\beta=0,$
further
$f(s)=\frac{1}{s}$
to get that
$\xi_{s}(x)=x+s.$
If instead of general one-parameter family of functions, we consider a sensitivity function $\xi$, then identity (3) expresses the fact that $\xi$ is the sensitivity function of an anchored and parallel psychometric family =$\mathcal{F}=\left\{p_{a}\,|\,a\in I\right\}$.
As the following statement shows this assumption is quite restrictive even without Iverson’s law of similarity. Below we describe those sensitivity functions that stem from an anchored, parallel, and balanced psychometric family.
Proposition 3.
Let =$\mathcal{F}=\left\{p_{a}\,|\,a\in I\right\}$ be a psychometric family that is anchored, parallel and balanced, and denote by $\xi$ the sensitivity function of $\mathcal{F}$. Then there exists a function $\nu\colon S\to\mathbb{R}$ which is antisymmetric with respect to the point $s_{0}=\frac{1}{2}$, i.e.
$\nu(1-s)=-\nu(s)\qquad\left(s\in S\right)$
such that
$\xi_{s}(x)=x+\nu(s)$
holds for all $x\in I$ and $s\in S$.
Proof.
Let =$\mathcal{F}=\left\{p_{a}\,|\,a\in I\right\}$ be a psychometric family that is anchored, parallel and balanced, and denote by $\xi$ the sensitivity function of $\mathcal{F}$. Due to the anchored and parallel properties of $\mathcal{F}$, there exist strictly increasing, continuous functions $u$ and $v$ such that
$\xi_{s}(x)=u(x)+v(s)\qquad\left(x\in I,s\in S\right)$
holds. Since $\mathcal{F}$ was assumed to be balanced as well, we have
$\xi_{1-s}\left(\xi_{s}(x)\right)=x$
for all possible $x\in I$ and $s\in S$. Using the above representation, this means that the functions $u$ and $v$ have to satisfy
$u(u(x)+v(s))=x-v(1-s)\qquad\left(x\in I,s\in S\right).$
The functions $u$ and $v$ are continuous and strictly increasing, thus the latter identity implies that
$u(x+s)=u^{-1}(x)-v(1-v^{-1}(s))\qquad\left(x\in I,s\in S\right),$
which is a Pexider equation. Thus, we have especially that
$u(x)=\alpha_{1}x+\beta_{1}\qquad\text{and}\qquad u^{-1}(x)=\alpha_{1}x+\beta_{2}\qquad\left(x\in I\right)$
with some appropriate constants $\alpha_{1},\beta_{1},\beta_{2}$. This is possible however only if $\alpha_{1}=1$ and $\beta_{2}=-\beta_{1}$. So
$v(s)+2\beta_{1}=-v(1-s)\qquad\left(s\in S\right).$
This means that the function $\nu\colon S\to\mathbb{R}$ defined by
$\nu(s)=v(s)+\beta_{1}\qquad\left(s\in S\right)$
is antisymmetric with respect to the point $s_{0}=\frac{1}{2}$. Therefore,
$\xi_{s}(x)=u(x)+v(s)=\left[x+\beta_{1}\right]+\left[\nu(s)-\beta_{1}\right]=x+\nu(s)$
for all $x\in I$ and $s\in S$, as stated. ∎
3.2 Multiplicatively translational mappings
In this subsection, we investigate Iverson’s law of similarity
$\xi_{s}(\lambda x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right)$
assuming that the two-variable mapping $\eta$ is multiplicatively translational.
Definition 1.
Let $S,J\subset\mathbb{R}$ be intervals such that $0\in S$ and $1\in J$. A mapping $\eta\colon J\times S\to S$ is called multiplicatively translational if
$\eta(\lambda\tilde{\lambda},s)=\eta(\tilde{\lambda},\eta(\lambda,s))$
and also the boundary conditions
$\eta(\lambda,0)=0\qquad\eta(1,s)=s$
hold for all $\lambda,\tilde{\lambda}\in J$ and $s\in S$.
A source of motivation for studying multiplicatively translational $\eta$ comes from the shift invariance relationship
$\xi_{\lambda^{\theta}s}(\lambda x)=\lambda\,\xi_{s}(x)$
appearing in the work of [Pavel and Iverson, 1981]. In this context, $\xi_{s}(x)$ represents the intensity of an unmasked tone that matches the loudness of a tone of intensity $x$ embedded in a background of intensity $s$. We could write this shift invariance relationship as
$\xi_{s}(\lambda x)=\lambda\,\xi_{\lambda^{-\theta}s}(x)\,,$
which is clearly a case of Iverson’s law of similarity with $\eta(\lambda,s)=\lambda^{-\theta}s$. Note that the multiplicatively translational property generalizes the transformation $\eta(\lambda,s)=\lambda^{-\theta}s$, since we have for this transformation that
$\eta(\tilde{\lambda},\eta(\lambda,s))=\eta(\tilde{\lambda},\lambda^{-\theta}s)=\tilde{\lambda}^{-\theta}\lambda^{-\theta}s=\eta(\lambda\tilde{\lambda},s)\,.$
We also point out that in this experimental context, it is reasonable that the shift invariance relationship would hold for $s=0$.
Note the following relationship between Iverson’s law of similarity and the multiplicatively translational property, also pointed out in [Hsu and Iverson, 2016]. If we write $\tilde{\lambda}x$ in place of $x$ in Iverson’s law of similarity, then by a successive application of the similarity law,
$\xi_{s}(\lambda\tilde{\lambda}x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(\tilde{\lambda}x)=\gamma(\lambda,s)\gamma(\tilde{\lambda},\eta(\lambda,s))\xi_{\eta(\tilde{\lambda},\eta(\lambda,s))}(x)\quad\left(\lambda,\tilde{\lambda}\in J,s\in S\right)$
follows. On the other hand, if we use Iverson’s law of similarity only once, but for $\lambda\tilde{\lambda}$, we obtain that
$\xi_{s}(\lambda\tilde{\lambda}x)=\gamma(\lambda\tilde{\lambda},s)\xi_{\eta(\lambda\tilde{\lambda},s)}(x)\quad\left(\lambda,\tilde{\lambda}\in J,s\in S\right).$
Thus necessarily,
$\gamma(\lambda,s)\gamma(\tilde{\lambda},\eta(\lambda,s))\xi_{\eta(\tilde{\lambda},\eta(\lambda,s))}(x)=\gamma(\lambda\tilde{\lambda},s)\xi_{\eta(\lambda\tilde{\lambda},s)}(x).\quad\left(\lambda,\tilde{\lambda}\in J,s\in S\right)$
Therefore, the assumption that $\eta$ is multiplicatively translational expresses in some sense a kind of consistency in the above computations.
Lemma 3.
Let $S,J\subset\mathbb{R}$ be intervals such that $0\in S$ and $1\in J$. If the mapping $\eta\colon J\times S\to S$ is multiplicatively translational and there exists an $s^{\ast}\in S$ such that the mapping
$J\ni\lambda\longmapsto\eta(\lambda,s^{\ast})$
is bijective, then there exists a bijective function $H\colon J\to S$ with $H(0)=0$ such that
$\eta(\lambda,s)=H(\lambda\cdot H^{-1}(s))\qquad\left(\lambda,s\in J\right).$
Proof.
Let $S,J\subset\mathbb{R}$ be intervals such that $0\in S$ and $1\in J$ and let $\eta\colon J\times S\to S$ be a multiplicatively translational mapping. Assume that there exists an $s^{\ast}\in S$ such that the mapping
$J\ni\lambda\longmapsto\eta(\lambda,s^{\ast}).$
is bijective. Define the function $H\colon J\to S$ through
$H(\lambda)=\eta(\lambda,s^{\ast})\qquad\left(\lambda\in J\right).$
Then $H$ is a bijection from $J$ onto $S$. Further, since $\eta$ is multiplicatively translational,
$H(\lambda\tilde{\lambda})=\eta(\lambda,H(\tilde{\lambda}))$
holds for all $\lambda,\tilde{\lambda}\in J$, from which
$\eta(\lambda,s)=H(\lambda\cdot H^{-1}(s))\qquad\left(\lambda\in J,s\in S\right)$
follows. Moreover, we also have
$0=\eta(\lambda,0)=H(\lambda\cdot H^{-1}(0))\qquad\left(\lambda\in J\right),$
from which we deduce
$H^{-1}(0)=\lambda H^{-1}(0)$
for all $\lambda\in J$, so $H^{-1}(0)=0$ and also $H(0)=0$. ∎
Remark 8.
Observe that from the previous proof, it follows that either $J=\left\{1\right\}$ and thus $S=\left\{0\right\}$ follows, or $0\in J$ holds. So $[0,1]\subset J$, otherwise the function $\eta$ is defined on the singleton $\left\{(1,0)\right\}$.
Remark 9.
We further note that in the above proof, the boundary conditions only play a role in that with their help we can show that $H(0)=0$. This means that if we omit these boundary conditions, the representation
$\eta(\lambda,s)=H(\lambda\cdot H^{-1}(s))$
is still true.
Remark 10.
Multiplicatively translational mappings were previously studied in [Hsu and Iverson, 2016] and even in [Aczél, 1966, Aczél, 1987]. However, we would like to emphasize that the conditions there are different from the ones we use. Indeed, according to [Aczél, 1987, Section 6, Theorem 3], if $S$ is an interval and $\eta\colon]0,+\infty[\times S\to S$ is a function that is continuous in both variables and for which the mapping
$\lambda\to\eta(\lambda,s)$
is nonconstant for every $s\in S$, and further
$\eta(\tilde{\lambda}\lambda,s)=\eta(\lambda,\eta(\lambda,s))$
holds for all $\tilde{\lambda},\lambda$ and $s\in S$, then there exists a continuous and strictly monotonic function $\psi\colon]0,+\infty[\to S$ such that
$\eta(\lambda,s)=\psi(\lambda\cdot\psi^{-1}(s))\qquad\left(\lambda\in]0,+\infty[,s\in S\right).$
Although the representations are the same, in this case it is assumed that $J=]0,+\infty[$ and the regularity conditions are different.
In what follows we will consider only those multiplicatively translational mappings $\eta$ for which there exists an $s^{\ast}\in S$ such that the mapping
$J\ni\lambda\longmapsto\eta(\lambda,s^{\ast})$
is bijective. In other words, we assume that there exists a bijective function $H\colon J\to S$ with $H(0)=0$ such that
$\eta(\lambda,s)=H(\lambda\cdot H^{-1}(s))\qquad\left(\lambda\in J,s\in S\right).$
For the sake of simplicity and easier distinction, such mappings will be termed regular multiplicatively translational mappings.
Observe that in this case the assumptions of Theorem 1 are fulfilled. Indeed, a careful adaption of that proof yields that we have
$\xi_{s^{\ast}}\left(\frac{H^{-1}(s)}{H^{-1}(s^{\ast})}x\right)=\gamma\left(\frac{H^{-1}(s)}{H^{-1}(s^{\ast})},s^{\ast}\right)\xi_{s}(x)\qquad\left(x\in I,s\in S\right).$
Proposition 4.
Let $I,J,S\subset\mathbb{R}$ be intervals with $I\cdot J=\left\{\lambda x\,|\,x\in I,\lambda\in J\right\}\subset I$ and $1\in J$, $0,1\in S$. Let further $\gamma,\eta\colon J\times S\to S$ be functions and assume that the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ fulfills
$\xi_{s}(\lambda x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right).$
If $\eta$ is a regular multiplicatively translational mapping then
$\gamma(\lambda,s)=\frac{\kappa(h(s)\lambda)}{\kappa(h(s))}\qquad\left(\lambda,s\in S\right)$
holds, except for the points where the denominator vanishes, where the functions $\kappa$ and $h$ are defined by
[…]
Proof.
Under the assumptions of the proposition, let $\tilde{\lambda}\in J$ be arbitrary and let us substitute $\tilde{\lambda}x$ in place of $x$ in Iverson’s law of similarity to get
$\xi_{s}(\lambda\tilde{\lambda}x)=\gamma(\lambda,s)\xi_{\eta(\lambda,s)}(\tilde{\lambda}x)=\gamma(\lambda,s)\gamma(\tilde{\lambda},\eta(\lambda,s))\xi_{\eta(\tilde{\lambda,\eta(\lambda,s)})}(x)\qquad\left(x\in I,\lambda,\tilde{\lambda}\in J,s\in S\right).$
On the other hand, Iverson’s law of similarity with $\lambda\tilde{\lambda}$ instead of $\lambda$ is
$\xi_{s}(\lambda\tilde{\lambda}x)=\gamma(\lambda\tilde{\lambda},s)\xi_{\eta(\lambda\tilde{\lambda},s)}(x)\qquad\left(x\in I,\lambda,\tilde{\lambda}\in J,s\in S\right).$
Using that the mapping $\eta$ is multiplicatively translational, we get that
$\gamma(\lambda\tilde{\lambda},s)=\gamma(\lambda,s)\cdot\gamma(\tilde{\lambda},\eta(\lambda,s))$
holds for all $\lambda,\tilde{\lambda}\in J,s\in S$. Further, since $\eta$ is not only multiplicatively translational but also regular multiplicatively translational, we have
$\gamma(\lambda\tilde{\lambda},s)=\gamma(\lambda,s)\cdot\gamma(\tilde{\lambda},H(\lambda\cdot H^{-1}(s)))\qquad\left(\lambda,\tilde{\lambda}\in J,s\in S\right).$
Let us substitute $s=1$ and define the functions $h,\kappa$ on $S$ by
$h(s)=\frac{H^{-1}(s)}{H^{-1}(1)}\qquad\text{and}\qquad\kappa(\lambda)=\gamma(\lambda,1)\qquad\left(\lambda\in J,s\in S\right)$
to deduce that
$\kappa(h(s)\lambda)=\kappa(h(s))\cdot\gamma(\lambda,s)\qquad\left(\lambda\in J,s\in S\right),$
that is,
$\gamma(\lambda,s)=\frac{\kappa(h(s)\lambda)}{\kappa(h(s))}$
holds for all $\lambda,s\in S$. ∎
Corollary 1.
Let $I,J,S\subset\mathbb{R}$ be intervals with $1\in I$, $0,1\in J$, $0,1\in S$, and $I\cdot J=\left\{\lambda x\,|\,x\in I,\lambda\in J\right\}\allowbreak\subset I$. Let further $\eta\colon J\times S\to S$ and $\phi\colon S\to\mathbb{R}$ be functions and assume that the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ fulfills the following power law
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right).$
If $\eta$ is a regular multiplicatively translational mapping, then there exists a function $F\colon I\to\mathbb{R}$ such that
$\xi_{s}(x)=x^{\phi(s)}F(x\cdot H^{-1}(s))\qquad\left(x\in I,s\in S\right).$
Proof.
Due to Proposition 4, we automatically have that
$\eta(\lambda,s)=H(\lambda\cdot H^{-1}(s))\qquad\text{and}\qquad\gamma(\lambda,s)=\lambda^{\phi(s)}\qquad\left(\lambda\in J,s\in S\right)$
with some function $H\colon J\to S$ such that $H(0)=0$. Further,
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{H(\lambda\cdot H^{-1}(s))}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right).$
With $x=1$, we get that
$\xi_{s}(\lambda)=\lambda^{\phi(s)}\xi_{H(\lambda\cdot H^{-1}(s))}(1)\qquad\left(x\in I,\lambda\in J,s\in S\right).$
Define the function $F\colon I\to\mathbb{R}$ by
$F(x)=\xi_{H(x)}(1)\qquad\left(x\in I\right)$
to get that
$\xi_{s}(x)=x^{\phi(s)}F(x\cdot H^{-1}(s))\qquad\left(x\in I,s\in S\right).$
In the statement below we consider the power law
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right)$
with a monotonic function $\phi$.
Corollary 2.
Under the assumptions of the previous corollary let us assume that the function $\phi\colon J\to J$ is monotonic on the interval $J$. Suppose the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in J)$ fulfills the following power law
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,\lambda,s\in J\right).$
Then there exists a countable partition of subintervals of $J$ such that $\phi$ is constant on each subinterval. That is, there exist $(J_{\alpha})_{\alpha\in A}$ such that $\bigcup_{\alpha\in A}J_{\alpha}=J$ and if $\alpha,\beta\in A$, $\alpha\neq\beta$, then $J_{\alpha}\cap J_{\beta}=\emptyset$, and real constants $c_{\alpha}$ for $\alpha\in A$ such that
$\phi(s)=c_{\alpha}\qquad\left(s\in J_{\alpha},\alpha\in A\right).$
In other words, the function $\phi$ is locally constant.
Proof.
Under the suppositions of the previous corollary, assume now that the function $\phi$ is monotonic on the interval $S$. Then the points of discontinuity of the function $\phi$ form a countable subset in $S$. Further this countable set induces a countable partition $(S_{\alpha})_{\alpha\in A}$ of the interval $S$ such that for all $\alpha\in A$, the function $\phi$ is continuous on $S_{\alpha}^{\circ}$ (on the interior of the interval $S_{\alpha}$).
Assume to the contrary that there exists an $\alpha\in A$ such that the interval $S_{\alpha}$ has a proper subinterval $\tilde{S}$ such that $\phi$ is strictly monotonic (and due to the construction, continuous) on $\tilde{S}$.
Due to Iverson’s law of similarity, we have
$\xi_{s}(\lambda\tilde{\lambda}x)=(\lambda\tilde{\lambda})^{\phi(s)}\xi_{\eta(\lambda\tilde{\lambda},s)}(x)\qquad\left(x\in I,\lambda,\tilde{\lambda}\in J,s\in\tilde{S}\right).$
On the other hand, we also have
$\xi_{s}(\lambda\tilde{\lambda}x)=\xi_{s}(\lambda(\tilde{\lambda x}))=\lambda^{\phi(s)}\cdot\xi_{\eta(\lambda,s)}(\tilde{\lambda}x)=\lambda^{\phi(s)}\cdot\tilde{\lambda}^{\phi(\eta(\lambda,s))}\xi_{\eta(\tilde{\lambda},\eta(\lambda,s))}(x)\qquad\left(x\in I,\lambda,\tilde{\lambda}\in J,s\in\tilde{S}\right).$
Since $\eta$ satisfies the regular multiplicative translation property, we have
$(\lambda\tilde{\lambda})^{\phi(s)}=\lambda^{\phi(s)}\cdot\tilde{\lambda}^{\phi(\eta(\lambda,s))},$
that is,
$\phi(s)=\phi(\eta(\lambda,s))$
for all $\lambda\in J,s\in\tilde{S}$. At the same time, the function $\phi$ was assumed to be continuous and strictly monotonic on $\tilde{S}$. Thus
$s=\eta(\lambda,s)$
holds for all $\lambda\in J,s\in\tilde{S}$. Since $\eta$ is a regular multiplicatively translational mapping,
$\eta(\lambda,s)=H(\lambda\cdot H^{-1}(s))$
holds for all $\lambda\in J,s\in S$ with an appropriate function $H$. This however means that we have
$s=H(\lambda\cdot H^{-1}(s)),$
i.e.,
$H^{-1}(s)=\lambda\cdot H^{-1}(s)$
for all $\lambda\in J,s\in\tilde{S}$ implying that $\lambda=1$ holds for all $\lambda\in J$. This is a contradiction. Thus $\phi$ cannot be strictly monotonic, but it is monotonic. ∎
Remark 11.
Notice that the assumption that the function $\eta$ is a regularly multiplicative mapping was used only at the end of the proof of Corollary 1. Indeed, to deduce that we have
$\phi(s)=\phi(\eta(\lambda,s)),$
we used only that $\eta$ fulfills $\eta(\lambda\tilde{\lambda},s)=\eta(\lambda,\eta(\tilde{\lambda},s))$. Thus, the proof of Corollary 2 shows that if $\eta$ is a function such that $\eta(\lambda\tilde{\lambda},s)=\eta(\lambda,\eta(\tilde{\lambda},s))$ is satisfied, then
  1. (a)
    either $\phi$ is locally constant (cf. Corollary 2)
  2. (b)
    or $\eta(\lambda,s)=s$ for all possible values $\lambda$ and $s$.
In the latter case we have Falmagne’s power law for the sensitivity functions $\xi_{s}\,(s\in S)$, i.e.,
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{s}(x)\qquad\left(x\in I,s\in S\right)$
holds. Recall that $\xi$ is termed to be weakly balanced if for all $x\in I$ we have $\xi_{\frac{1}{2}}(x)=x$. According to [Hsu and Iverson, 2016, Theorem 6], if, in addition to the above, $\xi$ admits a weakly balanced affine representation and $\phi$ is continuous, strictly monotonic and nowhere zero, then $\xi$ has a Fechnerian representation. The mentioned [Hsu and Iverson, 2016] paper also contains results about the power law
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,s\in S\right)$
with a multiplicatively translational mapping $\eta$. Recall that in the case where $\eta$ is regularly multiplacatively translational, the conditions of Theorem 1 hold. Thus, with the help of Proposition 1, we can specially obtain those sensitivities that satisfy the above power law and have an affine representation. In such a way we deduce [Hsu and Iverson, 2016, Corollary 7] without assuming continuous differentiability about the scales $u$ and $\sigma$ and twice differentiability about $\xi$.
Proposition 5.
Let $I,J\subset\mathbb{R}$ be intervals with $I\cdot J=\left\{\lambda x\,|\,x\in I,\lambda\in J\right\}\subset I$ and $0,1\in J$, $0\in S$. Let further $\eta\colon J\times S\to S$ be a regular multiplicatively translational mapping and $\phi\colon S\to\mathbb{R}$ be a monotonic function and assume that the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ fulfills the following power law
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right).$
Then the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ cannot admit a Fechnerian representation, i.e., there is no continuous and strictly monotonic function $u\colon I\to\mathbb{R}$ such that
$u(\xi_{s}(x))-u(x)=s\qquad\left(x\in I,s\in S\right).$
Proof.
Assume to the contrary that $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ do admit a Fechnerian representation, that is, let us assume that there exists a continuous and strictly monotonic function $u\colon I\to\mathbb{R}$ such that
$u(\xi_{s}(x))-u(x)=s\qquad\left(x\in I,s\in S\right).$
Using the former results, we get that the function $\phi$ must be locally constant. Thus there exists an interval $\tilde{S}\subset S$ containing the point $s=0$ such that
$\phi(s)=\rho\qquad\left(s\in\tilde{S}\right).$
We show that the above representation cannot hold on this interval $\tilde{S}$. If we show that the desired representation is impossible on this subinterval $\tilde{S}$, then it already follows that it is not fulfilled on the larger interval $S$ either.
Thus, from now on we will work on the interval $\tilde{S}$. Therefore for the sake of simplicity, we can assume that $\tilde{S}=S$. Then on this interval, the sensitivity functions $\xi_{s}\,(s\in S)$ are of the form
$\xi_{s}(x)=x^{\rho}F(x\cdot H^{-1}(s))\qquad\left(x\in I,s\in S\right).$
Writing this to the Fechnerian representation, we arrive to
$u\left(x^{\rho}F(x\cdot H^{-1}(s))\right)-u(x)=s\qquad\left(x\in I,s\in S\right).$
From this, with the substitution $s=0$,
$u(x^{\rho}\cdot F(0))-u(x)=0\qquad\left(x\in I\right)$
follows. Since the function $u$ is strictly monotonic, this is possible only if $\rho=1$ and $F(0)=1$. Therefore the above equation is reduced to
$u(x\cdot F(x\cdot H^{-1}(s)))=s+u(x)\qquad\left(x\in I,s\in S\right),$
or equivalently
$u(x\cdot F(xs))=H(s)+u(x)\qquad\left(x\in I,s\in S\right).$
Let $s\in S$, $s\neq 0$ be arbitrary and let us substitute $\dfrac{x}{s}$ in place of $x$ to deduce that
$u\left(\frac{x}{s}\cdot F(1)\right)=H(s)+u\left(\frac{x}{s}\right)\qquad\left(x\in I,s\in S,s\neq 0\right),$
i.e.,
$\tilde{u}\left(\frac{x}{s}\right)=H(s)$
holds for the function $\tilde{u}$ defined by
$\tilde{u}(x)=u(xF(1))-u(x)\qquad\left(x\in I\right).$
Let us now choose pair $(x,s),(x^{*},s^{*})\in I\times S\setminus\left\{0\right\}$ so that $s\neq s^{*}$ and $\dfrac{x}{s}=\dfrac{x^{*}}{s^{*}}$, then we have
$H(s)=\tilde{u}\left(\frac{x}{s}\right)=\tilde{u}\left(\frac{x^{*}}{s^{*}}\right)=H(s^{*}),$
contrary to the strict monotonicity of the function $H$. ∎
According to Corollary 1, if the one-parameter family of functions $\xi$ satisfies the power law
$\xi_{s}(\lambda x)=\lambda^{\phi(s)}\xi_{\eta(\lambda,s)}(x)$
with some regular multiplicatively translational mapping $\eta$, then $\xi$ can be written in the form
$\xi_{s}(x)=x^{\phi(s)}F(xH^{-1}(s)).$
Afterwards, we showed that if $\phi$ is monotone, then it is necessarily locally constant. So we have
$\xi_{s}(x)=x^{r}F(xH^{-1}(s)).$
In Proposition 5, we proved that then $\xi$ cannot admit a Fechnerian representation. This suggests that the assumptions in the above statements – although they seem natural – are collectively too restrictive.
In the following, we examine whether one-parameter families of the form
$\xi_{s}(x)=x^{r}F(xH^{-1}(s))$
can have a subtractive representation. As we shall see, the answer is yes. Note that one-parameter families of this form automatically satisfy the above power law, with the choice $\eta(\lambda,s)=H(\lambda\cdot H^{-1}(s))$ and $\phi(s)=r$. In this case $\eta$ is obviously multiplicatively translational, but not necessarily regular multiplicatively translational.
Proposition 6.
Let $I$ and $S$ be real intervals of positive length and suppose that the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ can be represented as
$\xi_{s}(x)=x^{r}F(x\cdot H^{-1}(s))\qquad\left(x\in I,s\in S\right)$ (4)
and also as
$\xi_{s}(x)=u^{-1}(s+w(x))\qquad\left(x\in I,s\in S\right)$ (5)
with an appropriate real constant $r$, continuous and strictly increasing functions $u,w$ and $H$ and with a continuous function $F$. If the domain of the functions $u,w$ and $H$ and the range of the function $F$ is contained in the set of positive reals, then the following cases are possible.
  1. Case I
    $\xi_{s}(x)=e^{\alpha}e^{sb}\cdot x^{r+1}\qquad\left(x\in I,s\in S\right)$
    with appropriate constants $\alpha$ and $b$.
  2. Case II
    $\xi_{s}(x)=e^{\alpha}(c+e^{-\delta s}(dx^{\kappa}-\beta c))^{\rho}\qquad\left(x\in I,s\in S\right)$
    with appropriate constants $\alpha,\beta,\delta,\kappa,\rho$ and $c$.
  3. Case III
    $\xi_{s}(x)=e^{\alpha}\left(cx^{\frac{r}{\rho}}+s-\varepsilon\right)^{\rho}\qquad\left(x\in I,s\in S\right)$
    with appropriate constants $\alpha,\varepsilon,\rho$ and $c$.
Proof.
The assumption that the domain of the functions $u,w$ and $H$ and the range of the function $F$ is contained in the set of positive reals guarantees that we can take the logarithm of both the sides of equation (4). In this case
$\log(\xi_{s}(x))$ $=r\log(x)+\log(F(x\cdot H^{-1}(s)))$
$=r\log(x)+\log\circ F\circ\exp(\log(x)+\log\circ H^{-1}(s))$
follows for all $x\in I$ and $s\in S$. Further, if we take the logarithm of both the sides of (5), we deduce
$\log(\xi_{s}(x))=\log\circ u^{-1}(s+w(x))\qquad\left(x\in I,s\in S\right).$
Therefore we have
$\log\circ u^{-1}(s+w(x))=r\cdot\log(x)+\log\circ F\circ\exp(\log(x)+\log\circ H^{-1}(s)),$
that is,
$\log\circ u^{-1}(s+y)=r\cdot\log\circ w^{-1}(y)+\log\circ F\circ\exp(\log\circ w^{-1}(y)+\log\circ H^{-1}(s))$
for all $x\in J$ and $s\in S$, where the interval $J$ denotes the range of the function $u$. Let us introduce the functions $h,m,f,\ell$ and $g$ by
$h$ $=\log\circ u^{-1}$
$m$ $=-r\cdot\log w^{-1}$
$f$ $=\log\circ F\circ\exp$
$\ell$ $=\log\circ w^{-1}$
$g$ $=\log\circ H^{-1}$
to obtain that the latter equation takes the form
$f(\ell(y)+g(s))=m(y)+h(y+s)\qquad\left(y\in J,s\in S\right).$
Since $h,\ell$ and $m$ are philandering, due to the respective properties of the functions $u$ and $w$, Lemma 2 can be applied. However, let us note that in our case with the above notations, we have $r\cdot l+m\equiv 0$. This shows that Cases IV and V are impossible in our case.
  1. Case I
    In this case
    $m(x)=\rho\ell(x)-\rho bx+\alpha+\rho\beta+\tau,$
    at the same time, due to $r\cdot l+m\equiv 0$,
    $\ell(x)=\frac{\rho b}{r+\rho}x-(\alpha+\rho\beta+\tau),$
    from which
    $w(x)=\frac{r+\rho}{\rho b}\log(x)+\frac{r+\rho}{\rho b}(\alpha+\rho\beta+\tau).$
    Further,
    $h(x)=-\tau+\rho bx,$
    so
    $u^{-1}(x)=e^{-\tau}e^{\rho bx}.$
    Therefore,
    $\xi_{s}(x)=e^{-\tau}e^{sb}\exp\left(\frac{r+\rho}{\rho b}(\alpha+\rho\beta+\tau)\right)\cdot x^{r+\rho}$
    holds. On the other hand,
    $g(x)=\beta+bx,$
    so
    $H^{-1}(s)=e^{\beta}e^{bs}$
    and
    $f(x)=\alpha+\rho x,$
    so
    $F(x)=e^{\alpha}x^{\rho}$
    implying that
    $\xi_{s}(x)=e^{\alpha}e^{\beta\rho}e^{b\rho s}x^{r+\rho}.$
    However, the two representations must agree in all points, so comparing them shows that either $b=0$ or $\rho=1$. From the first possibility, it would follow that the inverse of the function $u$ is a constant, which is impossible. So $\rho=1$. Thus we have
    $\xi_{s}(x)=e^{\alpha}e^{sb}\cdot x^{r+1}\qquad\left(x\in I,s\in S\right)$
    with appropriate constants $\alpha$ and $b$.
  2. Case II
    In this case
    $h(x)=-\tau+\alpha+\rho\log(bc+de^{\delta x}),$
    yielding that
    $u^{-1}(x)=e^{-\tau+\alpha}\left(bc+de^{\delta x}\right)^{\rho}$
    and
    $\ell(x)=-\frac{1}{\kappa}\log\left(\beta+be^{-\delta x}\right),$
    so
    $w(x)=-\frac{1}{\delta}\log\left(\frac{x^{-\kappa}-\beta}{b}\right).$
    From these, it follows that
    $\xi_{s}(x)=e^{-\tau+\alpha}\left(bc+de^{\delta s}\frac{b}{x^{-\kappa}-\beta}\right).$
    At the same time, since
    $g(x)=\frac{1}{\kappa}\log(-\beta c+de^{\delta x}),$
    we get that
    $H^{-1}(s)=\left(-\beta c+de^{\delta s}\right)^{\frac{1}{\kappa}}.$
    Further,
    $f(x)=\alpha+\rho\log(c+e^{\kappa x}),$
    we obtain that
    $F(x)=e^{\alpha}(c+x^{\kappa})^{\rho}.$
    Therefore
    $\xi_{s}(x)=x^{r}e^{\alpha}\left(c+x^{\kappa}(-\beta c+de^{\delta s})\right)^{\rho}.$
    Again, the obtained representations for $\xi$ have to agree everywhere, so
    $\xi_{s}(x)=e^{\alpha}(c+e^{-\delta s}(dx^{\kappa}-\beta c))^{\rho}$
    holds with some appropriate constants.
  3. Case III
    In case of the last possibility
    $h(x)=\rho\log(bd\alpha x+\alpha\varepsilon-b\beta)$
    so
    $u^{-1}(x)=\left(bd\alpha x+\alpha\varepsilon-b\right)^{\rho}$
    and since
    $\ell(x)=-\frac{1}{\kappa}\log(\varepsilon+bdx),$
    we have
    $w(x)=\frac{x^{\kappa}-\varepsilon}{bd}.$
    Therefore,
    $\xi_{s}(x)=\left(bd\alpha s+\alpha x^{-\kappa}-b\beta\right)^{\rho}.$
    At the same time, since
    $g(x)=\frac{1}{\kappa}\log\left(\beta-bdx\right),$
    we get that
    $H^{-1}(s)=\left(\beta-bds\right)^{\frac{1}{\kappa}}$
    and
    $f(x)=\rho\log\left(\alpha-be^{\kappa x}\right),$
    thus
    $F(x)=\left(\alpha-bx^{\kappa}\right)^{\rho}.$
    So
    $\xi_{s}(x)$ $=x^{r}\left(\alpha-bx^{\kappa}(\beta-bds)\right)^{\rho}$
    $=\left(\alpha x^{\frac{r}{\rho}}-bx^{\frac{r}{\rho}+\kappa}(\beta-bds)\right)^{\rho}$
    $=\left(\alpha x^{\frac{r}{\rho}}-b\beta+b^{2}ds\right)^{\rho},$
    where we also used that $\frac{r}{\rho}+\kappa=0$. The latter relationship results from the equation $r\cdot l+m\equiv 0$. Therefore,
    $\xi_{s}(x)=e^{\alpha}\left(cx^{\frac{r}{\rho}}+s-\varepsilon\right)^{\rho}$
    holds with some appropriate constants.
Regarding the assumptions of Proposition 6, note that the domain of the functions $u,w$ and $H$ and the range of the function $F$ will be contained in the set of positive reals if the sets $I$ and $J$ are contained in the set of positive reals and the outputs of $\xi_{s}$ are positive. These are reasonable assumptions in models of intensity discrimination.
Remark 12.
Observe that the mapping $\xi$ obtained in Case I corresponds to Case I of [Doble and Hsu, 2020, Theorem 10] if we choose there the function $\eta$ to be
$\eta(\lambda,s)=s+\frac{1}{b}\log(\lambda).$
Case II in the above theorem is the same as that of [Doble and Hsu, 2020, Theorem 10]. Finally, Case III agrees with Case V of [Doble and Hsu, 2020, Theorem 10]. The constants are denoted differently here than in the mentioned paper, but the mentioned representations are the same.
Corollary 3.
Let $I,J,S\subset\mathbb{R}$ be intervals with $I\cdot J=\left\{\lambda x\,|\,x\in I,\lambda\in J\right\}\subset I$ and $0,1\in J$, $0\in S$. Let further $\eta\colon J\times S\to S$ and $g\colon J\to\mathbb{R}$ be functions and assume that the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ fulfills
$\xi_{s}(\lambda x)=g(\lambda)\xi_{\eta(\lambda,s)}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right).$
If $\eta$ is a regular multiplicatively translational mapping and $g$ is measurable on some subinterval $\tilde{J}$ of $J$ with positive length, then there exist real numbers $\alpha$ and $\mu$ and a function $F\colon I\to\mathbb{R}$ such that
[…]
Proof.
Due to Proposition 4, we have
$\frac{\kappa(h(s)\lambda)}{\kappa(h(s))}=g(\lambda)\qquad\left(\lambda\in J,s\in S\right),$
that is,
$\kappa(h(s)\lambda)=\kappa(h(s))g(\lambda)\qquad\left(\lambda\in J,s\in S\right),$
or, in other words,
$\kappa(\lambda s)=g(\lambda)\kappa(s)\qquad\left(\lambda\in J,s\in S\right).$
This equation is a multiplicative Pexider equation. Further the function $g$ is measurable on some subinterval $\tilde{J}$ of $J$ with positive length. Thus $\kappa$ and $g$ are constant multiples of a power function. Therefore, there exist real numbers $\alpha$ and $\mu$ such that
$g(\lambda)=\alpha\lambda^{\mu}\qquad\left(\lambda\in J\right)$
and
$\xi_{s}(x)=\alpha x^{\mu}F(x\cdot H^{-1}(s))\qquad\left(x\in I,s\in S\right).$
Let $S,J\subset\mathbb{R}$ be intervals such that $0\in S$ and $1\in J$ holds. Let further $\theta$ be a fixed real number and let us consider the mapping $\eta\colon J\times S\to\mathbb{R}$ defined by
$\eta(\lambda,s)=\lambda^{\theta}s\qquad\left(\lambda\in J,s\in S\right).$
Recall that in this case
$\eta(\lambda\tilde{\lambda},s)=\eta(\tilde{\lambda},\eta(\lambda,s))\qquad\left(\lambda,\tilde{\lambda}\in J,s\in S\right)$
and
$\eta(1,s)=s\qquad\text{and}\qquad\eta(\lambda,0)=0\qquad\left(\lambda\in J,s\in S\right)$
hold. In other words, the above mapping $\eta$ is multiplicatively translational.
We should emphasize that in Lemma 3 we provided a formula for regular multiplicatively translational mappings, that is, for mappings that also have the property that there exists an $s^{\ast}\in S$ such that the mapping
$J\ni\lambda\longmapsto\eta(\lambda,s^{\ast})$
is bijective. This means that mappings of the above form, i.e.,
$\eta(\lambda,s)=\lambda^{\theta}s\qquad\left(\lambda\in J,s\in S\right)$
cannot necessarily be represented in the form
$\eta(\lambda,s)=H(\lambda\cdot H^{-1}(s))\qquad\left(\lambda\in J,s\in S\right).$
Indeed, consider the case when $\theta=-1$, i.e.,
$\eta(\lambda,s)=\frac{s}{\lambda}.$
Then for all fixed $s^{\ast}\in S$, the mapping
$J\ni\lambda\longmapsto\eta(\lambda,s^{\ast})=\frac{s^{\ast}}{\lambda}\in S$
cannot be bijective, since due to the boundary conditions we should have $0\in S$, but $\dfrac{s^{\ast}}{\lambda}\neq 0$ for all $\lambda\in J$.
Regardless, the cases when $\eta$ is not regular multiplicatively translational, but only multiplicatively translational, can be important from the point of view of applications, see [Pavel and Iverson, 1981] and [Hsu and Iverson, 2016].
We would also like to illustrate this case with an example below. Before all this, however, we examine the case when $\theta$ is positive in the above example.
Corollary 4.
Let $I,J,S\subset\mathbb{R}$ be intervals with $I\cdot J=\left\{\lambda x\,|\,x\in I,\lambda\in J\right\}\subset I$ and $0,1\in J$, $0\in S$. Let further $\theta>0$ be arbitrarily fixed, $\gamma\colon J\times S\to\mathbb{R}$ be a function and assume that the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ fulfills the following shift invariance
$\xi_{s}(\lambda x)=\gamma(\lambda,s)\xi_{\lambda^{\theta}s}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right).$
Then there exist functions $f,F\colon I\to\mathbb{R}$ such that
[…]
except for those points $s\in S$ for which we have $f(s^{\frac{1}{\theta}})=0$.
Proof.
Consider the function $H$ defined on $J$ by
$H(s)=s^{\theta}\qquad\left(s\in S\right).$
Then
$H(\lambda\cdot H^{-1}(s))=\lambda^{\theta}s$
holds for all $\lambda\in J,s\in S$. This means that Proposition 4 applies and the function $\gamma$ has to fulfill
$\gamma(\lambda,s)\cdot\gamma(\tilde{\lambda},\lambda^{\theta}s)=\gamma(\lambda\tilde{\lambda},s)$
for all $\lambda,\tilde{\lambda}\in J,s\in S$. If $s^{\ast}\in S\setminus\left\{0\right\}$ is arbitrarily fixed, then we get from the above identity with $s=s^{\ast}$ that
$\gamma(\lambda,s^{\ast})\cdot\gamma(\tilde{\lambda},\lambda^{\theta}s^{\ast})=\gamma(\lambda\tilde{\lambda},s^{\ast})\qquad\left(\lambda,\tilde{\lambda}\in J\right).$
Introducing the function
$f(\lambda)=\gamma({s^{\ast}}^{-\frac{1}{\theta}}\lambda,s^{\ast})\qquad\left(\lambda\in J\right),$
we get that
$\gamma(\lambda,s)=\frac{f(\lambda s^{\frac{1}{\theta}})}{f(s^{\frac{1}{\theta}})}$
holds for all $\lambda\in J,s\in S\setminus\left\{0\right\}$. Writing this representation back to the shift-invariance property, we get that
$\xi_{s}(\lambda x)=\frac{f(\lambda s^{\frac{1}{\theta}})}{f(s^{\frac{1}{\theta}})}\xi_{\lambda^{\theta}s}(x)$
for all $x\in I$ and $\lambda\in J,s\in S\setminus\left\{0\right\}$, yielding that
$\xi_{s}(x)=\frac{f(xs^{\frac{1}{\theta}})}{f(s^{\frac{1}{\theta}})}\cdot F(x^{\theta}s)\qquad\left(x\in I,s\in S\setminus\left\{0\right\}\right)$
with an appropriate function $F$. ∎
In the following statement, we will consider the special shift-invariance correcponding to $\theta=-1$ in the work [Iverson and Pavel, 1981]. Note that, due to the comments above, Corollary 4 cannot be applied in this case, since the mapping
$\eta(\lambda,s)=\frac{s}{\lambda}$
is not regular multiplicatively translational, even though it has the multiplicatively translational property. Accordingly, other ideas and tools are needed to investigate this case.
Proposition 7.
Let $S,I\subset\mathbb{R}$ and $J\subset]0,+\infty[$ be intervals such that $1\in J$, $0\in S$, $I\cdot J\subset I$ and $\left\{\frac{s}{\lambda}\,|\,s\in S,\lambda\in J\right\}\subset J$. If the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ fulfills
$\xi_{s}(\lambda x)=\lambda\xi_{\frac{s}{\lambda}}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right),$
then, there exists a real constant $c$ and a function $\varphi\colon\mathbb{R}\to\mathbb{R}$ such that
$\xi_{s}(x)=\begin{cases}cx,&\text{ if }\,s=0\\ s\cdot\varphi\left(\frac{x}{s}\right),&\text{ if }\,s\neq 0\end{cases}\qquad\left(x\in I,s\in S\right).$
Proof.
Let $S,I\subset\mathbb{R}$ and $J\subset]0,+\infty[$ be intervals such that $1\in J$, $0\in S$, $I\cdot J\subset I$ and $\left\{\frac{s}{\lambda}\,|\,s\in S,\lambda\in J\right\}\subset J$. Assume the one-parameter family of functions $\xi_{s}\colon I\to\mathbb{R}\,(s\in S)$ fulfills
$\xi_{s}(\lambda x)=\lambda\xi_{\frac{s}{\lambda}}(x)\qquad\left(x\in I,\lambda\in J,s\in S\right).$
With $s=0$ we obtain the multiplicative Pexider equation
$\xi_{0}(\lambda x)=\lambda\xi_{0}(x)\qquad\left(x\in I,\lambda\in J\right).$
Thus
$\xi_{0}(x)=cx\qquad\left(x\in I\right)$
with an appropriate real constant $c$. Further, with a fixed $s^{\ast}\neq 0$ we deduce that
$\xi_{s^{\ast}}(\lambda x)=\lambda\xi_{\frac{s^{\ast}}{\lambda}}(x)\qquad\left(x\in I,\lambda\in J\right).$
In other words,
$\xi_{s}(x)=\frac{s}{s^{\ast}}\xi_{s^{\ast}}\left(\frac{s^{\ast}}{s}x\right)$
holds for all $x\in I$ and $s\in S\setminus\left\{0\right\}$. Thus, if we consider the function $\varphi$ defined on $I$ by
$\varphi(x)=\frac{1}{s^{\ast}}\xi_{s^{\ast}}(s^{\ast}x)\qquad\left(x\in I\right),$
it can be seen that the statement of the proposition is fulfilled, indeed. ∎
In summary, in this subsection we examined the following. Suppose that there is a one-parameter family of functions $\xi$ that satisfies Iverson’s law of similarity. In addition, suppose that the function $\eta$ is regular multiplicative translational. The statements above tell us that in this case both the functions $\xi$ and $\gamma$ have some special form. The question of which of these one-parameter families $\xi$ satisfy some psychophysical representation will be answered in one of our future papers.
Finally, we also compare the current results with the results of [Doble and Hsu, 2020]. In that paper those sensitivity functions that satisfy Iverson’s law of similarity and also admit a subtractive representation were determined. Note that this automatically entails the fulfillment of certain regularity conditions, which we did not assume in this paper.
Nevertheless, it may be interesting to raise the question of which of the functions $\eta$ in [Doble and Hsu, 2020] have the regular multiplicatively translational property. As we will see below, the boundary conditions appearing in the definition of this notion are quite restrictive.
Indeed, an easy computation shows that all the functions (these appear in Cases II, IV, and V in [Doble and Hsu, 2020])
$\begin{array}[]{rcl}\eta_{1}(\lambda,s)&=&-\dfrac{1}{\kappa}\log\left(\lambda^{-\delta}(e^{-\kappa s}-\beta)+\beta\right)\\ \eta_{2}(\lambda,s)&=&s+\dfrac{\delta}{\kappa}\log(\lambda)\\ \eta_{3}(\lambda,s)&=&\lambda^{-\delta}(s-\varepsilon)+\varepsilon\end{array}$
fulfill equation
$\eta(\lambda\tilde{\lambda},s)=\eta(\lambda,\eta(\tilde{\lambda},s))\qquad\left(\lambda\in J,s\in J\right).$
However, the boundary conditions in the definition are satisfied only if
  • $\beta=1$ or $\delta=0$ in case of the function $\eta_{1}$, i.e.,
    $\eta_{1}(\lambda,s)=-\dfrac{1}{\kappa}\log\left(\frac{e^{-\kappa s}-1}{\lambda^{\delta}}+1\right)\quad\text{or}\quad\eta_{1}(\lambda,s)=s\qquad\left(\lambda\in J,s\in S\right).$
  • $\delta=0$ in case of the function $\eta_{2}$, i.e.,
    $\eta_{2}(\lambda,s)=s\qquad\left(\lambda\in J,s\in S\right).$
  • $\delta=0$ or $\varepsilon=0$ in case of the function $\eta_{3}$, i.e.,
    $\eta_{3}(\lambda,s)=s\quad\text{or}\quad\eta_{3}(\lambda,s)=s\lambda^{-\delta}\qquad\left(\lambda\in J,s\in S\right).$
Note that these are the multiplicatively translational mappings, of which the function
$\eta(\lambda,s)=s$
is not regular. In this case, however, Iverson’s law of similarity is a multiplicative type equation. Thus if we assume e.g. that for all fixed $s\in S$ is monotonic on $I$, then we obtain that
$\xi_{s}(x)=c(s)x^{\varphi(s)}\qquad\left(x\in I,s\in S\right)$
with some appropriate functions $c,\varphi\colon S\to\mathbb{R}$.
Further, the mapping
$\eta(\lambda,s)=s\cdot\lambda^{-\delta}$
is regular multiplicatively translational and this case corresponds to the shift-invariance dealt with in Corollary 4.
Finally, if we consider only regular multiplicatively translational mappings, then in the first case of $\eta_{1}$ we have
$H(s)=-\dfrac{1}{\kappa}\log\left(\frac{\alpha}{s^{\delta}}+1\right)\qquad\left(s\in S\right).$
Since now we have $[0,1]\subset J$, the parameter $\delta$ must be negative. So we may write $-\delta$ in place of $\delta$ (with some positive $\delta$). Further we set $\kappa=-\log(\alpha+1)$, which guarantees $H(1)=1$. The latter is only a technical condition that makes calculations easier. This brings us to the last case, i.e., when
$H(s)=\dfrac{\log(\alpha s^{\delta}+1)}{\log(\alpha+1)}\qquad\left(s\in S\right),$
in other words,
$\eta(\lambda,s)=\dfrac{\log\left(\lambda^{\delta}\left(e^{\log(\alpha+1)s}-1\right)+1\right)}{\log(\alpha+1)}\qquad\left(\lambda\in J,s\in S\right).$
Note that in this case, we have
$H^{-1}(s)=\left(\frac{\exp(\log(\alpha+1)s)-1}{\alpha}\right)^{\frac{1}{\rho}}\qquad\left(s\in S\right).$
Given the argument presented at the beginning of this section,
$\xi_{s^{\ast}}\left(\frac{H^{-1}(s)}{H^{-1}(s^{\ast})}x\right)=\gamma\left(\frac{H^{-1}(s)}{H^{-1}(s^{\ast})},s^{\ast}\right)\xi_{s}(x)\qquad\left(x\in I,s\in S\right)$
with some fixed $s^{\ast}\in S$. Assume now that there exists $\bar{s}\in S$ and a subinterval $\bar{S}\subset S$ such that $\gamma(\lambda,\bar{s})\neq 0$ for all $\lambda\in J$. Then
$\xi_{s}(x)=\dfrac{\Phi(H^{-1}(s)x)}{g(H^{-1}(s))}\qquad\left(x\in I,s\in S\right)$
holds with some appropriate functions $\Phi$ and $g$. The question of which one-parameter functions of this form satisfy a psychometric representation may also be answered in a future paper.